### v1

# Soliton solutions to the Sawada–Kotera equation

Abstract
We consider the direct and inverse scattering problems for the third-order differential equation in the reflectionless case. We formulate a corresponding Riemann–Hilbert problem using input consisting of the bound-state poles of a transmission coefficient and the bound-state dependency constants. With the time-evolved dependency constants, using the solution to the Riemann–Hilbert problem, we construct soliton solutions to an integrable system of fifth-order nonlinear partial differential equations. By imposing some appropriate restrictions on the dependency constants, we show that those soliton solutions yield soliton solutions to the Sawada–Kotera equation.
AMS Subject Classification (2020): 34A55 34M50 35C08
Keywords: inverse scattering for the third-order equation, bound-state dependency constants, soliton solutions, Sawada–Kotera equation
1 Introduction
In this paper we are interested in soliton solutions to an integrable system of coupled fifth-order nonlinear partial differential equations and, in particular, one of its special cases, namely, the Sawada–Kotera equation [26]. We construct such solutions by solving the inverse scattering problem for a third-order differential equation in the reflectionless case. The solution to the inverse scattering problem is obtained by solving a related Riemann–Hilbert problem using input consisting of the bound-state poles of a transmission coefficient and the bound-state dependency constants. Using the time-evolved dependency constants in the input data set, we show that the solution to the inverse problem yields soliton solutions to the aforementioned integrable nonlinear system and also soliton solutions to the Sawada–Kotera equation. This method explains the physical origin of the $2\mathbf{N}$ real parameters appearing in the relevant $\mathbf{N}$-soliton solution formula obtained by the bilinear method [17] of Hirota, by relating $\mathbf{N}$ of the real parameters to the bound-state energy values and relating the remaining $\mathbf{N}$ real parameters to the bound-state dependency constants.
Our paper presents the construction of soliton solutions to the Sawada–Kotera equation via the inverse scattering transform method [14]. Such solutions to the Sawada–Kotera equation are usually obtained by using Hirota’s bilinear method or a slight modification [16, 24] of that method. Hirota’s bilinear method is an effective algebraic method to obtain soliton solutions to various integrable nonlinear partial differential equations. However, it is an ad hoc procedure, and it does not provide any insight or physical motivation for the construction of those soliton solutions. In particular, it does not explain how the parameters appearing in those soliton solutions may be related to any physical quantities. On the other hand, the method we use to construct those soliton solutions is a fundamental procedure that can be applied to a wide variety of other integrable evolution equations, and it also relates the parameters appearing in soliton solutions to the bound-state poles and bound-state dependency constants for the relevant ordinary linear differential operator.
In the inverse scattering transform method, a time-evolved potential is constructed from the time-evolved scattering data. A soliton solution corresponds to a time-evolved potential corresponding to a reflectionless scattering data set. A reflectionless scattering data set comprises the bound-state information only. The bound-state information in turn consists of the bound-state energies and the bound-state normalization constants, or equivalently it consists of the bound-state poles of a transmission coefficient and a bound-state dependency constant for each bound-state pole. For example, an $\mathbf{N}$-soliton solution to the Sawada–Kotera equation is a real-valued function of $x$ and $t$ and it contains $2\mathbf{N}$ real parameters. As our method indicates, $\mathbf{N}$ of those real parameters identify the location of the bound-state poles of a transmission coefficient and the remaining $\mathbf{N}$ real parameters each identify the bound-state dependency constants associated with the bound-state poles.
Our paper is organized as follows. In Section 2 we introduce the integrable system (2.6) of coupled nonlinear partial differential equations associated with the third-order linear equation (2.8). We indicate how the Sawada–Kotera equation arises as a special case of (2.6) by uncoupling the integrable system (2.6). In Section 3 we provide a summary of the direct scattering problem for (2.8) in the reflectionless case. This is done by introducing the relevant solutions to (2.8) at each $k$-value in the complex $k$-plane. We indicate how the left and right transmission coefficients are related to the spacial asymptotics of two relevant solution to (2.8). In Section 4 we introduce the bound-state poles of the left transmission coefficient $T_{\text{\rm{l}}}(k),$ and we describe the bound-state dependency constant for each bound-state pole. For simplicity we assume that each bound-state pole is simple. In Section 5 we describe the solution to the inverse scattering problem for (2.8) in the reflectionless case. This is done by providing some explicit expressions for the potentials $Q$ and $P$ in terms of the input data set consisting of $T_{\text{\rm{l}}}(k),$ the bound-state poles of $T_{\text{\rm{l}}}(k),$ and the bound-state dependency constants specified for the bound-state poles. In the case of the recovery of the potentials $Q$ and $P$ that depend on the parameter $t,$ the use of time-evolved dependency constants yields soliton solutions to the corresponding integrable system. In Section 6, by imposing the appropriate restrictions on a certain solution to the nonlinear system (2.6), we obtain the $\mathbf{N}$-soliton solution to the Sawada–Kotera equation (2.1). Finally, in Section 7, we illustrate the $\mathbf{N}$-soliton solution to the Sawada–Kotera equation when $\mathbf{N}=1,2,3.$
2 The integrable system and the Sawada–Kotera equation
The Sawada–Kotera equation is the fifth-order nonlinear partial differential equation given by
$Q_{t}+Q_{xxxxx}+5\,Q_{x}\,Q_{xx}+5\,Q\,Q_{xxx}+5\,Q^{2}\,Q_{x}=0,\qquad x,\,t\in\mathbb{R},$ (2.1)
where the subscripts denote the corresponding partial derivatives. It is used as a model to describe the propagation of surface water waves in long, narrow, shallow canals, similar to the modeling used by the KdV (Korteweg–de Vries) equation [19]. For both the KdV equation and the Sawada–Kotera equation, we assume that the solutions are real valued. Hence, the quantity $Q$ appearing in (2.1) is assumed to be real valued. Compared to third-order nonlinear KdV equation, the Sawada–Kotera equation takes into account higher-order nonlinear and dispersive effects. In Section 7 of our paper, we remark on some different behaviors of soliton solutions to the Sawada–Kotera equation and to the KdV equation.
The Sawada–Kotera equation is integrable in the sense of the inverse scattering transform. This is due to the fact that it has a Lax pair, i.e. it is related to two linear differential operators $L$ and $A$ satisfying the Lax operator equation [20]
$L_{t}+LA-AL=0,$ (2.2)
where $L_{t}$ is obtained by applying the partial $t$-derivative on $L.$ By using the corresponding Lax pair $(L,A)$ on the left-hand side of (2.2), that left-hand side becomes the zero multiplication operator when (2.1) holds. There are actually two different Lax pairs for the Sawada–Kotera equation (2.1). The first pair $(L,A)$ is given by
$\begin{cases}L=D^{3}+QD,\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr A=9\,D^{5}+15\,QD^{3}+15\,Q_{x}\,D^{2}+\big{(}10\,Q_{xx}+5\,Q^{2}\big{)}D,\end{cases}$ (2.3)
where we have let $D:=d/dx$ and $D^{n}:=d^{n}/dx^{n}$ for $n\geq 2.$ The second Lax pair $(L,A)$ for (2.1) is given by
$\begin{cases}L=D^{3}+QD+Q_{x},\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr A=9\,D^{5}+15\,QD^{3}+30\,Q_{x}\,D^{2}+\big{(}25\,Q_{xx}+5\,Q^{2}\big{)}D+\big{(}10\,Q_{xxx}+10\,Q\,Q_{x}\big{)}.\end{cases}$ (2.4)
The Lax pairs in (2.3) and (2.4) can be obtained as the two special cases of the general Lax pair $(L,A)$ given by
$\begin{cases}L=D^{3}+QD+P,\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr A=9\,D^{5}+15\,QD^{3}+\big{(}15\,P+15\,Q_{x}\big{)}D^{2}+\big{(}15\,P_{x}+10\,Q_{xx}+5\,Q^{2}\big{)}D+\big{(}10\,P_{xx}+10\,Q\,P\big{)},\end{cases}$ (2.5)
where the potentials $Q$ and $P$ are independent of each other and are allowed to take complex values. The Lax pair in (2.3) is obtained from (2.5) when $Q$ is real valued and we have $P\equiv 0,$ and the Lax pair in (2.4) is obtained when $Q$ is real valued and we have $P=Q_{x}$. The general Lax pair $(L,A)$ given in (2.5) satisfies the Lax operator equation (2.2) provided the potentials $Q$ and $P$ appearing in (2.5) satisfy the fifth-order coupled nonlinear partial differential equations
$\begin{cases}Q_{t}+Q_{xxxxx}+5\,Q_{x}\,Q_{xx}+5\,Q\,Q_{xxx}+5\,Q^{2}\,Q_{x}+15\,Q_{xx}\,P+15\,Q_{x}\,P_{x}-30\,P\,P_{x}=0,\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr\begin{aligned} P_{t}+P_{xxxxx}+5\,Q\,P_{xxx}+15\,Q_{x}\,P_{xx}+20\,Q_{xx}\,P_{x}+&5\,Q^{2}\,P_{x}+10\,Q_{xxx}\,P-15\,P\,P_{xx}\\ &+10\,Q\,Q_{x}P-15\,(P_{x})^{2}=0.\end{aligned}\end{cases}$ (2.6)
In order to uncouple the nonlinear system (2.6), from the last three terms of the first coupled equation in (2.6) we see that we must have
$15\,Q_{xx}\,P+15\,Q_{x}\,P_{x}-30\,P\,P_{x}=0,$
which is equivalent to
$\left[(Q_{x}-P)\,P\right]_{x}=0.$ (2.7)
Integrating both sides of (2.7) with respect to $x$ and using the fact that $Q$ and $P$ must vanish as $x\to\pm\infty$ for each fixed $t,$ we obtain
$(Q_{x}-P)\,P=0,$
which shows that the uncoupling in (2.6) occurs when $P\equiv 0$ or when $P=Q_{x}.$ For each of those two cases, one can directly verify that the left-hand side of the second coupled equation in (2.6) vanishes identically. Hence, we confirm that the use of $P\equiv 0$ or $P=Q_{x}$ in (2.6) yields (2.1).
Associated with the linear operator $L$ in (2.5), we have the third-order ordinary linear differential equation
$\psi^{\prime\prime\prime}+Q(x)\,\psi^{\prime}+P(x)\,\psi=k^{3}\,\psi,\qquad x\in\mathbb{R}.$ (2.8)
We suppress the dependence on $t$ for the potentials $Q$ and $P,$ and we write $Q(x)$ and $P(x),$ respectively, instead of $Q(x,t)$ and $P(x,t).$ The adjoint equation [3] for (2.8) is given by
$\overline{\psi}^{\prime\prime\prime}+\overline{Q}(x)\,\overline{\psi}^{\prime}+\overline{P}(x)\,\overline{\psi}=k^{3}\,\overline{\psi},\qquad x\in\mathbb{R},$ (2.9)
with the adjoint potentials $\overline{Q}$ and $\overline{P}$ related to $Q$ and $P$ as
$\overline{Q}(x):=Q(x)^{\ast},\qquad\overline{P}(x):=Q^{\prime}(x)^{\ast}-P(x)^{\ast},\qquad x\in\mathbb{R},$ (2.10)
where we use an asterisk for complex conjugation and use an overbar to denote the quantities associated with the adjoint equation. Comparing (2.8) and (2.9), with the help of (2.10) we observe that, when the potential $Q$ is real valued and the potential $P$ is zero, the adjoint potential $\overline{Q}$ is equal to $Q$ and the adjoint potential $\overline{P}$ becomes equal to $Q_{x}.$ In that case, the equation in (2.8) reduces to
$\psi^{\prime\prime\prime}+Q(x)\,\psi^{\prime}=k^{3}\,\psi,\qquad x\in\mathbb{R},$ (2.11)
and the adjoint equation in (2.9) reduces to
$\overline{\psi}^{\prime\prime\prime}+Q(x)\,\overline{\psi}^{\prime}+Q_{x}(x)\,\overline{\psi}=k^{3}\,\overline{\psi},\qquad x\in\mathbb{R}.$ (2.12)
The lax pair in (2.3) is associated with the reduced equation (2.11) and the lax pair in (2.4) is associated with the reduced adjoint equation (2.12). This explains why the Sawada–Kotera equation (2.1) has two distinct Lax pairs given in (2.3) and (2.4), respectively.
3 The direct scattering problem in the reflectionless case
In this section we present the basic ingredients for the direct scattering problem for (2.8) in the reflectionless case. We assume that the two complex-valued potentials $Q$ and $P$ appearing in (2.8) each belong to the Schwartz class in $x\in\mathbb{R}$ for each fixed value of $t.$ That particular direct problem consists of the determination of the solutions to (2.8) and the left and right transmission coefficients $T_{\text{\rm{l}}}(k)$ and $T_{\text{\rm{r}}}(k),$ respectively, for (2.8) when the potentials $Q$ and $P$ are known. For the description of the direct scattering problem when the reflection coefficients are not zero, we refer the reader to [3, 18, 27].
It is convenient to divide the complex $k$-plane into four open sectors $\Omega_{1},$ $\Omega_{2},$ $\Omega_{3},$ $\Omega_{4}$ as indicated on the left plot of Figure 3.1 by using the directed half lines $\mathcal{L}_{1},$ $\mathcal{L}_{2},$ $\mathcal{L}_{3},$ $\mathcal{L}_{4},$ which are parametrized as
$\mathcal{L}_{1}:=\{k\in\mathbb{C}:k=zs\text{\rm{ for }}s\in[0,+\infty)\},$ (3.1)
$\mathcal{L}_{2}:=\{k\in\mathbb{C}:k=z^{2}s\text{\rm{ for }}s\in[0,+\infty)\},$
$\mathcal{L}_{3}:=\{k\in\mathbb{C}:k=-zs\text{\rm{ for }}s\in[0,+\infty)\},$ (3.2)
$\mathcal{L}_{4}:=\{k\in\mathbb{C}:k=-z^{2}s\text{\rm{ for }}s\in[0,+\infty)\}.$
We use $z$ to denote the special complex number $e^{2\pi i/3},$ which is also expressed as
$z:=-\displaystyle\frac{1}{2}+i\,\displaystyle\frac{\sqrt{3}}{2}.$ (3.3)
The open sectors $\Omega_{1},$ $\Omega_{2},$ $\Omega_{3},$ $\Omega_{4}$ are described by using the parametrizations given by
$\Omega_{1}:=\left\{k\in\mathbb{C}:\displaystyle\frac{2\pi}{3}<\arg[k]<\displaystyle\frac{4\pi}{3}\right\},$
$\Omega_{2}:=\left\{k\in\mathbb{C}:-\displaystyle\frac{2\pi}{3}<\arg[k]<-\displaystyle\frac{\pi}{3}\right\},$
$\Omega_{3}:=\left\{k\in\mathbb{C}:-\displaystyle\frac{\pi}{3}<\arg[k]<\displaystyle\frac{\pi}{3}\right\},$
$\Omega_{4}:=\left\{k\in\mathbb{C}:\displaystyle\frac{\pi}{3}<\arg[k]<\displaystyle\frac{2\pi}{3}\right\},$
where $\arg[k]$ denotes the argument function taking values in the interval $(-2\pi/3,4\pi/3).$ We use $\overline{\Omega_{1}},$ $\overline{\Omega_{2}},$ $\overline{\Omega_{3}},$ $\overline{\Omega_{4}}$ to denote the closures of the open sectors $\Omega_{1},$ $\Omega_{2},$ $\Omega_{3},$ $\Omega_{4},$ respectively, where we recall that each closure is obtained by adding the boundaries to the corresponding open sector.
We define the left Jost solution $f(k,x)$ and the right Jost solution $g(k,x)$ as the solutions to (2.8) with the respective spacial asymptotics given by
$\begin{cases}f(k,x)=e^{kx}\left[1+o(1)\right],\qquad x\to+\infty,\quad k\in\overline{\Omega_{1}},\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr f^{\prime}(k,x)=k\,e^{kx}\left[1+o(1)\right],\qquad x\to+\infty,\quad k\in\overline{\Omega_{1}},\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr f^{\prime\prime}(k,x)=k^{2}\,e^{kx}\left[1+o(1)\right],\qquad x\to+\infty,\quad k\in\overline{\Omega_{1}},\end{cases}$ (3.4)
$\begin{cases}g(k,x)=e^{kx}\left[1+o(1)\right],\qquad x\to-\infty,\quad k\in\overline{\Omega_{3}},\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr g^{\prime}(k,x)=k\,e^{kx}\left[1+o(1)\right],\qquad x\to-\infty,\quad k\in\overline{\Omega_{3}},\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr g^{\prime\prime}(k,x)=k^{2}\,e^{kx}\left[1+o(1)\right],\qquad x\to-\infty,\quad k\in\overline{\Omega_{3}}.\end{cases}$ (3.5)
When the reflection coefficients for (2.8) are all zero, the left transmission coefficient $T_{\text{\rm{l}}}(k)$ and the right transmission coefficient $T_{\text{\rm{r}}}(k)$ are obtained by using the appropriate respective spacial asymptotics of the Jost solutions $f(k,x)$ and $g(k,x).$ We have
$f(k,x)=e^{kx}T_{\text{\rm{l}}}(k)^{-1}\left[1+o(1)\right],\qquad x\to-\infty,\quad k\in\overline{\Omega_{1}},$
$g(k,x)=e^{kx}T_{\text{\rm{r}}}(k)^{-1}\left[1+o(1)\right],\qquad x\to+\infty,\quad k\in\overline{\Omega_{3}}.$
When the potentials $Q$ and $P$ in (2.8) belong to the Schwartz class in $x\in\mathbb{R},$ the transmission coefficients $T_{\text{\rm{l}}}(k)$ and $T_{\text{\rm{r}}}(k)$ have meromorphic extensions from their respective $k$-domains $\overline{\Omega_{1}}$ and $\overline{\Omega_{3}}$ to the entire complex $k$-plane. In fact, those extensions satisfy [3]
$T_{\text{\rm{r}}}(k)=\displaystyle\frac{1}{T_{\text{\rm{l}}}(k)},\qquad k\in\mathbb{C}.$
It is known [3] that (2.8) has the particular solutions $m(k,x)$ and $n(k,x)$ with the respective $k$-domains $\overline{\Omega_{2}}$ and $\overline{\Omega_{3}},$ and in the reflectionless case they satisfy the spacial asymptotics given by
$m(k,x)=e^{kx}\left[1+o(1)\right],\qquad x\to-\infty,\quad k\in\overline{\Omega_{2}},$
$n(k,x)=e^{kx}\left[1+o(1)\right],\qquad x\to-\infty,\quad k\in\overline{\Omega_{4}}.$
In the reflectionless case, we have [3] the spacial asymptotics
$m(k,x)=e^{kx}\,T_{\text{\rm{l}}}(z^{2}k)^{-1}\,T_{\text{\rm{r}}}(zk)[1+o(1)],\qquad x\to+\infty,\quad k\in\overline{\Omega_{2}},$
$n(k,x)=e^{kx}\,T_{\text{\rm{l}}}(zk)^{-1}\,T_{\text{\rm{r}}}(z^{2}k)[1+o(1)],\qquad x\to+\infty,\quad k\in\overline{\Omega_{4}},$
where we recall that $z$ is the cube root of unity appearing in (3.3).
Since $k$ appears as $k^{3}$ in (2.8), it follows that $\psi(zk,x)$ and $\psi(z^{2}k,x)$ are both solutions to (2.8) whenever $\psi(k,x)$ is a solution. The $k$-domain of $\psi(zk,x)$ is obtained from the $k$-domain of $\psi(k,x)$ by a clockwise rotation of $2\pi/3$ radians around the origin of the complex $k$-plane. Similarly, the $k$-domain of $\psi(z^{2}k,x)$ is obtained from the $k$-domain of $\psi(k,x)$ by a clockwise rotation of $4\pi/3$ radians. This allows us to determine three linearly independent solutions to (2.8) at each $k$-value in the complex $k$-plane. Toward that goal, we divide $\Omega_{1}$ into the open subsectors $\Omega_{1}^{\text{\rm{up}}}$ and $\Omega_{1}^{\text{\rm{down}}}$ by using the negative real axis $\mathbb{R}^{-},$ and we divide the open sector $\Omega_{3}$ into the open subsectors $\Omega_{3}^{\text{\rm{down}}}$ and $\Omega_{3}^{\text{\rm{up}}}$ by using the positive real axis $\mathbb{R}^{+}.$ On the middle plot of Figure 3.1, we show the partitioning of the complex $k$-plane into the six open sectors $\Omega_{1}^{\text{\rm{up}}},$ $\Omega_{1}^{\text{\rm{down}}},$ $\Omega_{2},$ $\Omega_{3}^{\text{\rm{down}}},$ $\Omega_{3}^{\text{\rm{up}}},$ and $\Omega_{4}.$ We use -$\mathbb{R}^{-}$ and $\mathbb{R}^{+},$ respectively, to denote the directed half lines parametrized as
$\mathbb{R}^{+}:=\{k=s:s\in[0,+\infty)\},$
$\mathbb{R}^{-}:=\{k=-s:s\in[0,+\infty)\}.$
On the right plot of Figure 3.1 we indicate three linearly independent solutions to (2.8) in each of the six closed sectors $\overline{\Omega_{1}^{\text{\rm{up}}}},$ $\overline{\Omega_{1}^{\text{\rm{down}}}},$ $\overline{\Omega_{2}},$ $\overline{\Omega_{3}^{\text{\rm{down}}}},$ $\overline{\Omega_{3}^{\text{\rm{up}}}},$ and $\overline{\Omega_{4}},$ where we recall that the overbar denotes the closure of the corresponding open sector.
Figure 3.1: The directed half lines $\mathcal{L}_{1},$ $\mathcal{L}_{2},$ $\mathcal{L}_{3},$ $\mathcal{L}_{4}$ and the open sectors $\Omega_{1},$ $\Omega_{2},$ $\Omega_{3},$ $\Omega_{4}$ in the complex $k$-plane are shown on the left plot. The complex $k$-plane is divided into the six sectors $\Omega^{\text{\rm{up}}}_{1},$ $\Omega^{\text{\rm{down}}}_{1},$ $\Omega_{2},$ $\Omega^{\text{\rm{down}}}_{3},$ $\Omega^{\text{\rm{up}}}_{3},$ and $\Omega_{4}$ as shown on the middle plot, and the $k$-domains of three basic solutions to (2.8) in each of the six regions, respectively, are shown on the right plot.
For each fixed real values of $x$ and $t,$ the large $k$-asymptotics of the basic solutions $f(k,x),$ $g(k,x),$ $m(k,x),$ and $n(k,x)$ are, respectively, given by [3]
$f(k,x)=e^{kx}\left[1+\displaystyle\frac{u_{1}(x)}{k}+\displaystyle\frac{u_{2}(x)}{k^{2}}+O\left(\displaystyle\frac{1}{k^{3}}\right)\right],\qquad k\to\infty\text{\rm{ in }}\overline{\Omega_{1}},$ (3.6)
$g(k,x)=e^{kx}\left[1+\displaystyle\frac{v_{1}(x)}{k}+\displaystyle\frac{v_{2}(x)}{k^{2}}+O\left(\displaystyle\frac{1}{k^{3}}\right)\right],\qquad k\to\infty\text{\rm{ in }}\overline{\Omega_{3}},$ (3.7)
$m(k,x)=e^{kx}\left[1+O\left(\displaystyle\frac{1}{k}\right)\right],\qquad k\to\infty\text{\rm{ in }}\overline{\Omega_{2}},$ (3.8)
$n(k,x)=e^{kx}\left[1+O\left(\displaystyle\frac{1}{k}\right)\right],\qquad k\to\infty\text{\rm{ in }}\overline{\Omega_{4}},$ (3.9)
where we have defined
$u_{1}(x):=\displaystyle\frac{1}{3}\int_{x}^{\infty}dy\,Q(y),\qquad x\in\mathbb{R},$ (3.10)
$u_{2}(x):=-\displaystyle\frac{1}{3}\int_{x}^{\infty}dy\left[Q^{\prime}(y)-P(y)\right]+\displaystyle\frac{1}{18}\left[\int_{x}^{\infty}dy\,Q(y)\right]^{2},\qquad x\in\mathbb{R},$ (3.11)
$v_{1}(x):=-\displaystyle\frac{1}{3}\int_{-\infty}^{x}dy\,Q(y),\qquad x\in\mathbb{R},$ (3.12)
$v_{2}(x):=\displaystyle\frac{1}{3}\int_{-\infty}^{x}dy\left[Q^{\prime}(y)-P(y)\right]+\displaystyle\frac{1}{18}\left[\int_{-\infty}^{x}dy\,Q(y)\right]^{2},\qquad x\in\mathbb{R}.$ (3.13)
Using (3.10)–(3.13), we express the potentials $Q$ and $P$ in terms of $u_{1}(x),$ $u_{2}(x),$ $v_{1}(x),$ $v_{2}(x)$ as
$Q(x)=-3\,\displaystyle\frac{du_{1}(x)}{dx},\qquad x\in\mathbb{R},$ (3.14)
$P(x)=3\left[u_{1}(x)\,\displaystyle\frac{du_{1}(x)}{dx}-\displaystyle\frac{d^{2}u_{1}(x)}{dx^{2}}-\displaystyle\frac{du_{2}(x)}{dx}\right],\qquad x\in\mathbb{R},$ (3.15)
$Q(x)=-3\,\displaystyle\frac{dv_{1}(x)}{dx},\qquad x\in\mathbb{R},$ (3.16)
$P(x)=3\left[v_{1}(x)\,\displaystyle\frac{dv_{1}(x)}{dx}-\displaystyle\frac{d^{2}v_{1}(x)}{dx^{2}}-\displaystyle\frac{dv_{2}(x)}{dx}\right],\qquad x\in\mathbb{R}.$ (3.17)
4 The bound states and the bound-state dependency constants
A bound state corresponds to a nontrivial solution to (2.8) which is square integrable in $x\in\mathbb{R}.$ If a bound state occurs at $k=k_{j}$ in the complex $k$-plane, then the number of linearly independent square-integrable solutions to (2.8) at $k=k_{j}$ determines the multiplicity of that bound state. In our paper, we only consider simple bound states, where the multiplicity of each bound state is equal to $1.$ We refer the reader to [4, 5, 6] for the treatment of bound states with multiplicities for various differential or difference equations. The analysis of solutions to (2.8) at $k=0$ is more challenging. When we analyze bound states for (2.8), since we assume that the potentials $Q$ and $P$ belong to the Schwartz class in $x\in\mathbb{R},$ we exclude the possibility of a bound state at $k=0.$
The bound states occur at the zeros of $T_{\text{\rm{l}}}(k)^{-1}$ in $\Omega_{1}^{\text{\rm{down}}}$ or $\Omega_{1}^{\text{\rm{up}}}.$ Without loss of generality, it is enough to consider the bound states at the poles of $T_{\text{\rm{l}}}(k)$ in $\Omega_{1}^{\text{\rm{down}}}.$ It turns out [3] that the analysis of the bound state at $k=k_{j}$ with $k_{j}\in\Omega_{1}^{\text{\rm{down}}}$ yields useful information also about the bound state at $k=k_{j}^{\ast}$ with $k_{j}^{\ast}\in\Omega_{1}^{\text{\rm{up}}}.$ We recall that we use an asterisk to denote complex conjugation.
We define the $3$-Wronskian of three functions $F(x),$ $G(x),$ and $H(x)$ as
$\left[F(x);G(x);H(x)\right]:=\begin{vmatrix}F(x)&G(x)&H(x)\\ F^{\prime}(x)&G^{\prime}(x)&H^{\prime}(x)\\ F^{\prime\prime}(x)&G^{\prime\prime}(x)&H^{\prime\prime}(x)\end{vmatrix}$,
where we have the determinant of the relevant $3\times 3$ matrix on the right-hand side. The $3$-Wronskian of any three solutions to (2.8) at any particular $k$-value is zero if and only if those three solutions are linearly dependent. Furthermore, because of the absence of $\psi^{\prime\prime}$ in (2.8), the $3$-Wronskian of those three solutions is independent of $x$ and its value can be evaluated at any particular $x$-value. For example, using their asymptotics as $x\to\pm\infty,$ we evaluate the $3$-Wronskian of $f(k,x),$ $g(zk,x),$ and $n(z^{2}k,x)$ in $\overline{\Omega_{1}^{\text{\rm{down}}}}$ as
$\left[f(k,x);g(zk,x);n(z^{2}k,x)\right]=-3z(1-z)k^{3}\,T_{\text{\rm{l}}}(k)^{-1},\qquad k\in\overline{\Omega_{1}^{\text{\rm{down}}}}.$ (4.1)
Let us assume that the nonzero complex constant $k_{j}$ is located in the open sector $\Omega_{1}^{\text{\rm{down}}}$ and that it corresponds to a bound state. Then, we have $T_{\text{\rm{l}}}(k_{j})^{-1}=0.$ From (4.1) we see that the three solutions $f(k_{j},x),$ $g(zk_{j},x),$ and $n(z^{2}k_{j},x)$ to (2.8) become linearly dependent. This allows us [3] to express $f(k_{j},x)$ as a linear combination of $g(zk_{j},x)$ and $n(z^{2}k_{j},x)$ as
$f(k_{j},x)=D(k_{j})\,g(zk_{j},x)+W(k_{j})\,n(z^{2}k_{j},x),\qquad x\in\mathbb{R},$ (4.2)
for some complex constants $D(k_{j})$ and $W(k_{j}).$
Let us divide the open sector $\Omega_{1}^{\text{\rm{down}}},$ shown on the middle plot of Figure 3.1, into two parts, the first of which is the sector with $\arg[k]\in(\pi,7\pi/6)$ and the second sector is described via $\arg[k]\in[7\pi/6,4\pi/3).$ The analysis in Section 3 of [3] shows that $D(k_{j})=0$ when $\arg[k_{j}]\in(\pi,7\pi/6),$ and hence (4.2) yields
$f(k_{j},x)=W(k_{j})\,n(z^{2}k_{j},x),\qquad\arg[k_{j}]\in\left(\pi,\displaystyle\frac{7\pi}{6}\right),\quad x\in\mathbb{R},$ (4.3)
where the nonzero complex constant $W(k_{j})$ corresponds to the dependency constant at the bound state with $k=k_{j}.$ On the other hand, from the analysis in Section 3 of [3] it follows that $W(k_{j})=0$ if we have $\arg[k_{j}]\in[7\pi/6,4\pi/3).$ In that case, from (4.2) with $W(k_{j})=0$ we get
$f(k_{j},x)=D(k_{j})\,g(zk_{j},x),\qquad\arg[k_{j}]\in\left[\displaystyle\frac{7\pi}{6},\displaystyle\frac{4\pi}{3}\right),\quad x\in\mathbb{R},$ (4.4)
with the nonzero complex constant $D(k_{j})$ corresponding to the dependency constant at the bound state with $k=k_{j}.$ We remark that, if the potentials $Q$ and $P$ in (2.8) depend also on the parameter $t,$ then the solutions to (2.8) also depend on $t.$ Consequently, the dependency constants $W(k_{j})$ and $D(k_{j})$ appearing in (4.3) and (4.4) also depend on $t.$ For further analysis related to the bound states of (2.8), we refer the reader to [3].
Let us consider the case where the potentials $Q$ and $P$ depend on the time variable $t$ with the time evolution governed by the linear operator $A$ given in the second line of (2.5). In that case, the bound-state dependency constants evolve in time as
$D(k_{j})=E(k_{j})\,e^{9(z^{2}-1)\,k_{j}^{5}\,t},\qquad\arg[k_{j}]\in\left[\displaystyle\frac{7\pi}{6},\displaystyle\frac{4\pi}{3}\right),$ (4.5)
$W(k_{j})=U(k_{j})\,e^{9(z-1)k_{j}^{5}t},\qquad\arg[k_{j}]\in\left(\pi,\displaystyle\frac{7\pi}{6}\right),$
where $E(k_{j})$ and $U(k_{j})$ are used to denote the values of the dependency constants $D(k_{j})$ and $W(k_{j}),$ respectively, at $t=0.$ We refer the reader to [3] for further information on the time evolution of the bound-state dependency constants.
5 The inverse scattering problem in the reflectionless case
In this section we present the basic results for the inverse problem for (2.8) in the reflectionless case. In that case, the inverse problem consists of the determination of the potentials $Q$ and $P$ when the transmission coefficients $T_{\text{\rm{l}}}(k)$ and $T_{\text{\rm{r}}}(k)$ are known. In general, the potentials $Q$ and $P$ cannot be determined uniquely when we only know the transmission coefficients. For the unique recovery of the potentials, we need some additional information. The specification of a dependency constant or a normalization constant at each bound state is a conventional way to recover the potentials uniquely. A bound-state dependency constant and a bound-state normalization constant are related [3] to each other via the residue of a transmission coefficient at the bound-state pole. The use of bound-state dependency constants is convenient when the potentials are recovered from the solution to a related Riemann-Hilbert problem. The use of bound-state normalization constants is appropriate when the potentials are recovered from the solution to a related linear integral equation, such as a Marchenko integral equation [7, 8, 10, 13, 21, 22] or a Gel’fand–Levitan integral equation [8, 21, 22].
In the reflectionless case, we can formulate the inverse problem for (2.8) as a Riemann–Hilbert problem in the complex $k$-plane. Thus, the transmission coefficients and the bound-state dependency constants are used as input to the Riemann-Hilbert problem, and the potentials are recovered from the solution to the Riemann-Hilbert problem. We refer the reader to Section 4 of [3] for the details of the analysis of the formulation of the corresponding Riemann–Hilbert problem. In this section, we present the relevant Riemann–Hilbert problem when a specific left transmission coefficient is used as input. The particular choice of the left transmission coefficient is motivated by the goal of obtaining soliton solutions to the Sawada–Kotera equation (2.1). We recall that we consider the reflectionless case for (2.8).
We choose the left transmission coefficient $T_{\text{\rm{l}}}(k)$ as
$T_{\text{\rm{l}}}(k)=\displaystyle\frac{\Gamma(k)}{\Gamma(-k)},$ (5.1)
where $\Gamma(k)$ is given by
$\Gamma(k):=\displaystyle\prod_{j=1}^{\mathbf{N}}(k+k_{j})(k+k_{j}^{\ast}).$ (5.2)
From (5.1) and (5.2) we see that the left transmission coefficient $T_{\text{\rm{l}}}(k)$ is defined on the whole complex $k$-plane and that it has $2{\mathbf{N}}$ poles located at $k=k_{j}$ and $k=k_{j}^{\ast},$ respectively, for $1\leq j\leq{\mathbf{N}}.$ The large $k$-asymptotics of $T_{\text{\rm{l}}}(k)$ is given by
$T_{\text{\rm{l}}}(k)=1+\displaystyle\frac{2\Sigma_{\mathbf{N}}}{k}+\displaystyle\frac{2\Sigma_{\mathbf{N}}^{2}}{k^{2}}+O\left(\displaystyle\frac{1}{k^{3}}\right),\qquad k\to\infty\text{\rm{ in }}\mathbb{C},$
where we have defined
$\Sigma_{\mathbf{N}}:=\sum_{j=1}^{\mathbf{N}}\left(k_{j}+k_{j}^{\ast}\right).$ (5.3)
The quantity $Q$ appearing in the Sawada–Kotera equation (2.1) is real valued. Since we would like to relate (5.1) eventually to (2.1), we are interested in the consideration of the potentials $Q$ and $P$ appearing in (2.8) only when they are real valued. Since the original $k$-domain of $T_{\text{\rm{l}}}(k)$ is the sector $\overline{\Omega_{1}},$ the restriction of $Q$ to the real values requires [3] that each of the $k_{j}$-values in (5.2) must occur on the directed half line with $\arg[k]=7\pi/6.$ Thus, we have a restriction on the locations of the bound-state poles. Since we only consider simple bound states, the restriction can be presented as
$k_{j}=iz\eta_{j},\qquad 1\leq j\leq{\mathbf{N}},$ (5.4)
where we recall that $z$ is the special complex constant in (3.3). Without loss of generality, we assume that $0<\eta_{1}<\dots<\eta_{\mathbf{N}}.$
We refer the reader to Section 4 of [3] for the formulation of the Riemann–Hilbert problem in the presence of reflection coefficients. Here, we present the basic steps of the formulation of the Riemann-Hilbert problem in the reflectionless case and we also describe the basic steps to solve that Riemann–Hilbert problem. For the details of the steps outlined here, we refer the reader also to Section 4 of [3]. Using the basic solutions $f(k,x),$ $g(k,x),$ $m(k,x),$ and $n(k,x)$ introduced in Section 3, we form the solutions $\Phi_{+}(k,x)$ and $\Phi_{-}(k,x)$ to (2.8) as
$\Phi_{+}(k,x):=\begin{cases}T_{\text{\rm{l}}}(k)\,f(k,x),\qquad k\in\overline{\Omega_{1}},\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr m(k,x),\qquad k\in\overline{\Omega_{2}},\end{cases}$ (5.5)
$\Phi_{-}(k,x):=\begin{cases}g(k,x),\qquad k\in\overline{\Omega_{3}},\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr n(k,x),\qquad k\in\overline{\Omega_{4}}.\end{cases}$ (5.6)
Using the directed half lines $\mathcal{L}_{1}$ and $\mathcal{L}_{3}$ defined in (3.1) and (3.2), respectively, we obtain the directed full line $\mathcal{L}$ via $\mathcal{L}:=\mathcal{L}_{1}\cup(-\mathcal{L}_{3}),$ where we recall that $\mathcal{L}_{3}$ is obtained from $\mathcal{L}_{3}$ by changing its direction. The parametrization of $\mathcal{L}$ is given by
$\mathcal{L}:=\{k\in\mathbb{C}:k=zs\text{\rm{ for }}s\in(-\infty,+\infty)\}.$
The directed line $\mathcal{L}$ divides the complex $k$-plane into two half planes $\mathcal{P}^{+}$ and $\mathcal{P}^{-}$ as shown in Figure 5.1.
Figure 5.1: The plus and minus regions in the complex $k$-plane are separated by the directed full line $\mathcal{L},$ and the plus and minus functions have their respective $k$-domains as indicated.
The open left-half complex plane $\mathcal{P}^{+}$ and the open right-half complex plane $\mathcal{P}^{-}$ can be parameterized as
$\mathcal{P}^{+}:=\{k\in\mathbb{C}:k=zs\text{\rm{ for }}s\in\mathbb{C}^{+}\},$
$\mathcal{P}^{-}:=\{k\in\mathbb{C}:k=zs\text{\rm{ for }}s\in\mathbb{C}^{-}\},$
where we use $\mathbb{C}^{+}$ and $\mathbb{C}^{-}$ to denote the upper-half and lower-half complex planes. We refer to $\mathcal{P}^{+}$ as the plus region and refer to $\mathcal{P}^{-}$ as the minus region, as indicated in Figure 5.1. We use $\overline{\mathcal{P}^{+}}$ and $\overline{\mathcal{P}^{-}}$ to denote their closures. Hence, we have $\overline{\mathcal{P}^{+}}:=\mathcal{P}^{+}\cup\mathcal{L}$ and $\overline{\mathcal{P}^{-}}:=\mathcal{P}^{-}\cup\mathcal{L}.$
We remark that the quantities $\Phi_{+}(k,x)$ and $\Phi_{-}(k,x)$ in (5.5) and (5.6), respectively, are certain specific solutions to (2.8). We proceed to set up our Riemann–Hilbert problem. It is known [3] in the reflectionless case that $\Phi_{+}(k,x)$ and $\Phi_{-}(k,x)$ satisfy
$\Phi_{+}(k,x)=\Phi_{-}(k,x),\qquad k\in\mathcal{L}.$ (5.7)
When (5.1), (5.2), and (5.4) hold, it is known [3] that, for each fixed $x\in\mathbb{R},$ the quantity $\Phi_{+}(k,x)$ is meromorphic in $k\in\mathcal{P}^{+}$ with simple poles located at $k=k_{j}$ and $k=k_{j}^{\ast}.$ On the other hand, for each fixed $x\in\mathbb{R},$ the quantity $\Phi_{-}(k,x)$ is analytic in $k\in\mathcal{P}^{-}.$ From (5.7) we conclude that $\Phi_{+}(k,x)$ is the meromorphic continuation of $\Phi_{-}(k,x)$ from $k\in\mathcal{P}^{-}$ to $k\in\mathcal{P}^{+},$ and that $\Phi_{-}(k,x)$ is the analytic continuation of $\Phi_{+}(k,x)$ from $k\in\mathcal{P}^{+}$ to $k\in\mathcal{P}^{-}.$ Multiplying both sides of (5.7) by $e^{-kx}\,\Gamma(-k),$ we get
$e^{-kx}\,\Gamma(-k)\,\Phi_{+}(k,x)=e^{-kx}\,\Gamma(-k)\,\Phi_{-}(k,x),\qquad k\in\mathcal{L}.$ (5.8)
For each fixed $x\in\mathbb{R},$ the left-hand side of (5.8) is analytic in $k\in\mathcal{P}^{+}$ and the right-hand side is analytic in $k\in\mathcal{P}^{-}.$ Hence, those two sides are analytic continuations of each other. Consequently, each side of (5.8) is entire with their respective analytic continuations in $k\in\mathbb{C}.$ With the help of (3.6)–(3.14), by using the generalized Liouville theorem [25], we conclude that each side of (5.8) is equal to a monic polynomial in $k$ of degree $2\mathbf{N},$ where the coefficients may depend on $x$ and $t.$ By suppressing the $t$-dependence in our notation, we obtain the solution to our Riemann–Hilbert problem posed in (5.7) as
$\Phi_{+}(k,x)=e^{kx}\,\displaystyle\frac{k^{2\mathbf{N}}+V(k)\,\mathbf{A}(x)}{\Gamma(-k)},\qquad k\in\overline{\mathcal{P}^{+}},\quad x\in\mathbb{R},$ (5.9)
$\Phi_{-}(k,x)=e^{kx}\,\displaystyle\frac{k^{2\mathbf{N}}+V(k)\,\mathbf{A}(x)}{\Gamma(-k)},\qquad k\in\overline{\mathcal{P}^{-}},\quad x\in\mathbb{R},$ (5.10)
where $V(k)$ is the row vector with $2\mathbf{N}$ components defined as
$V(k):=\begin{bmatrix}k^{2\mathbf{N}-1}&k^{2\mathbf{N}-2}&\cdots&k&1\end{bmatrix},\qquad k\in\mathbb{C},$
and $\mathbf{A}(x)$ is a column vector with $2\mathbf{N}$ entries that are functions of $x$ and $t.$ Because of the aforementioned meromorphic extensions, each of (5.9) and (5.10) holds when $k\in\mathbb{C}.$ We write $\mathbf{A}(x)$ in terms of its components as
$\mathbf{A}(x)=\begin{bmatrix}A_{2\mathbf{N}-1}(x)\\ A_{2\mathbf{N}-2}(x)\\ \vdots\\ A_{1}(x)\\ A_{0}(x)\end{bmatrix},\qquad x\in\mathbb{R}.$
We emphasize that we suppress the $t$-dependence in our notation for $\mathbf{A}(x)$ and its components $A_{j}(x)$ for $0\leq j\leq 2\mathbf{N}-1.$
We remark that the solutions $\Phi_{+}(k,x)$ and $\Phi_{-}(k,x)$ appearing in (5.9) and (5.10) comprise the general solution to the Riemann–Hilbert problem (5.7). In other words, those solutions contain $2\mathbf{N}$ arbitrary constants not depending on $k$ but depending on $x$ and $t$ through the scalar functions $A_{j}(x)$ for $0\leq j\leq 2\mathbf{N}-1.$ From (5.9), with the help of (5.1) and the first line of (5.5), we get
$f(k,x)=e^{kx}\,\displaystyle\frac{k^{2\mathbf{N}}+V(k)\,\mathbf{A}(x)}{\Gamma(k)},\qquad k\in\overline{\Omega_{1}},\quad x\in\mathbb{R}.$ (5.11)
From (5.9) and the second line of (5.5), we get
$m(k,x)=e^{kx}\,\displaystyle\frac{k^{2\mathbf{N}}+V(k)\,\mathbf{A}(x)}{\Gamma(-k)},\qquad k\in\overline{\Omega_{2}},\quad x\in\mathbb{R}.$ (5.12)
Similarly, from (5.10), with the help of (5.6), we have
$g(k,x)=e^{kx}\,\displaystyle\frac{k^{2\mathbf{N}}+V(k)\,\mathbf{A}(x)}{\Gamma(-k)},\qquad k\in\overline{\Omega_{3}},\quad x\in\mathbb{R},$ (5.13)
$n(k,x)=e^{kx}\,\displaystyle\frac{k^{2\mathbf{N}}+V(k)\,\mathbf{A}(x)}{\Gamma(-k)},\qquad k\in\overline{\Omega_{4}},\quad x\in\mathbb{R}.$ (5.14)
With the help of (5.2), we observe that the right-hand sides in (5.12)–(5.14) are analytic in their respective $k$-domains and that the right-hand side in (5.11) is meromorphic in its $k$-domain. Thus, we have confirmed the aforementioned meromorphic and analytic properties stated below (5.7).
In order to get a unique solution to (5.7), we use the dependency constants $D(k_{j})$ and $D(k_{j}^{\ast})$ at the bound states at $k=k_{j}$ and $k=k_{j}^{\ast},$ respectively, as follows. We recall that $D(k_{j})$ appears in (4.4) and (4.5). It is known [3] that when $k_{j}$ has the form given in (5.4), the bound-state dependency constants $D(k_{j})$ and $D(k_{j}^{\ast})$ can be written in simpler forms by introducing the modified bound-state dependency constant $\gamma_{j}$ by letting
$\gamma_{j}:=-\displaystyle\frac{\Gamma(k_{j})\,E(k_{j})}{\Gamma(-zk_{j})},\qquad 1\leq j\leq\mathbf{N},$ (5.15)
where we recall that $E(k_{j})$ and $\Gamma(k)$ are the quantities appearing in (4.5) and (5.2), respectively. To use the bound-state dependency constants $D(k_{j})$ and $D(k_{j}^{\ast})$ in the solution to the Riemann–Hilbert problem (5.7), we proceed as follows. With the help of (4.5), (5.11), and (5.13), from (4.4) we obtain
$e^{k_{j}x}\,\displaystyle\frac{k_{j}^{2\mathbf{N}}+V(k_{j})\,\mathbf{A}(x)}{\Gamma(k_{j})}=E(k_{j})\,\,e^{9(z^{2}-1)k_{j}^{5}t}e^{zk_{j}x}\,\displaystyle\frac{(zk_{j})^{2\mathbf{N}}+V(zk_{j})\,\mathbf{A}(x)}{\Gamma(-zk_{j})},\qquad 1\leq j\leq\mathbf{N}.$ (5.16)
The analog of (4.4) at $k=k_{j}^{\ast}$ is given by [3]
$f(k_{j}^{\ast},x)=D(k_{j}^{\ast})\,g(z^{2}k_{j}^{\ast},x),\qquad\arg[k_{j}^{\ast}]\in\left(\displaystyle\frac{2\pi}{3},\displaystyle\frac{5\pi}{6}\right],$ (5.17)
and the analog of (4.5) at $k=k_{j}^{\ast}$ is given by [3]
$D(k_{j}^{\ast})=E(k_{j}^{\ast})\,e^{9(z-1)\,(k_{j}^{\ast})^{5}\,t},\qquad\arg[k_{j}^{\ast}]\in\left(\displaystyle\frac{2\pi}{3},\displaystyle\frac{5\pi}{6}\right].$ (5.18)
With the help of (5.11), (5.13), and (5.18), from (5.17) we get
$\begin{split}e^{k_{j}^{\ast}x}\,&\displaystyle\frac{(k_{j}^{\ast})^{2\mathbf{N}}+V(k_{j}^{\ast})\,\mathbf{A}(x)}{\Gamma(k_{j}^{\ast})}\\ &\phantom{xxxxx}=E(k_{j}^{\ast})\,e^{9(z-1)(k_{j}^{\ast})^{5}t}\,e^{z^{2}k_{j}^{\ast}x}\,\displaystyle\frac{(z^{2}k_{j}^{\ast})^{2\mathbf{N}}+V(z^{2}k_{j}^{\ast})\,\mathbf{A}(x)}{\Gamma(-z^{2}k_{j}^{\ast})},\qquad 1\leq j\leq\mathbf{N}.\end{split}$ (5.19)
Let us introduce the quantity $\chi_{j},$ which is an exponential function of $x$ and $t,$ as
$\chi_{j}:=e^{(z-1)k_{j}x+9(z^{2}-1)k_{j}^{5}t},\qquad 1\leq j\leq\mathbf{N}.$ (5.20)
When $k_{j}$ has the form given in (5.4), from (5.20) we conclude that
$\chi_{j}^{\ast}=\chi_{j},\qquad 1\leq j\leq\mathbf{N},$
which indicates that each $\chi_{j}$ is real valued. Similarly, when $k_{j}$ is as in (5.4), with the help of (4.5) and (5.2) where the quantities $\Gamma(k)$ and $E(k_{j})$ appear, respectively, we get
$\frac{\Gamma(k_{j}^{\ast})\,E(k_{j}^{\ast})}{\Gamma(-z^{2}k_{j}^{\ast})}=\left(\displaystyle\frac{\Gamma(k_{j})\,E(k_{j})}{\Gamma(-zk_{j})}\right)^{\ast},\qquad 1\leq j\leq\mathbf{N}.$ (5.21)
Then, by comparing (5.21) with the right-hand side of (5.15), we conclude that the modified bound-state dependency constant at $k=k_{j}^{\ast}$ is equal to $\gamma_{j}^{\ast},$ which is the complex conjugate of the modified bound-state dependency constant $\gamma_{j}$ at $k=k_{j}$ appearing in (5.15). Consequently, we can rewrite (5.16) and (5.19), respectively, as
$k_{j}^{2\mathbf{N}}+V(k_{j})\,\mathbf{A}(x)=-\gamma_{j}\,\chi_{j}\left[(zk_{j})^{2\mathbf{N}}+V(zk_{j})\,\mathbf{A}(x)\right],\qquad 1\leq j\leq\mathbf{N},$ (5.22)
$(k_{j}^{\ast})^{2\mathbf{N}}+V(k_{j}^{\ast})\,\mathbf{A}(x)=-\gamma_{j}^{\ast}\,\chi_{j}\left[(z^{2}k_{j}^{\ast})^{2\mathbf{N}}+V(z^{2}k_{j}^{\ast})\,\mathbf{A}(x)\right],\qquad 1\leq j\leq\mathbf{N}.$ (5.23)
We remark that (5.23) comprises a linear algebraic system of $2\mathbf{N}$ equations for the $2\mathbf{N}$ unknowns $A_{j}(x)$ with $0\leq j\leq 2\mathbf{N}-1.$ We write that linear algebraic system in the matrix form as
$\mathbf{M}(x)\,\mathbf{A}(x)=-\mathbf{B}(x),$ (5.24)
where we have defined the $2\mathbf{N}\times 2\mathbf{N}$ matrix $\mathbf{M}(x)$ and the column vector $\mathbf{B}(x)$ with $2\mathbf{N}$ components as
$\mathbf{M}(x):=\begin{bmatrix}m_{2\mathbf{N}-1}(k_{1})&m_{2\mathbf{N}-2}(k_{1})&\cdots&m_{1}(k_{1})&m_{0}(k_{1})\\ m_{2\mathbf{N}-1}(k_{1}^{\ast})&m_{2\mathbf{N}-2}(k_{1}^{\ast})&\cdots&m_{1}(k_{1}^{\ast})&m_{0}(k_{1}^{\ast})\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ m_{2\mathbf{N}-1}(k_{\mathbf{N}})&m_{2\mathbf{N}-2}(k_{\mathbf{N}})&\cdots&m_{1}(k_{\mathbf{N}})&m_{0}(k_{\mathbf{N}})\\ m_{2\mathbf{N}-1}(k_{\mathbf{N}}^{\ast})&m_{2\mathbf{N}-2}(k_{\mathbf{N}}^{\ast})&\cdots&m_{1}(k_{\mathbf{N}}^{\ast})&m_{0}(k_{\mathbf{N}}^{\ast})\end{bmatrix}$, (5.25)
$\mathbf{B}(x):=\begin{bmatrix}m_{2\mathbf{N}}(k_{1})\\ m_{2\mathbf{N}}(k_{1}^{\ast})\\ \vdots\\ m_{2\mathbf{N}}(k_{\mathbf{N}})\\ m_{2\mathbf{N}}(k_{\mathbf{N}}^{\ast})\end{bmatrix}$. (5.26)
The entries appearing in (5.25) and (5.26) are given as
$\begin{cases}m_{l}(k_{j}):=k_{j}^{l}+(zk_{j})^{l}\,\gamma_{j}\,\chi_{j},\qquad 1\leq j\leq\mathbf{N},\quad 0\leq l\leq 2\mathbf{N},\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr m_{l}(k_{j}^{\ast}):=(k_{j}^{\ast})^{l}+(z^{2}k_{j}^{\ast})^{l}\,\gamma_{j}^{\ast}\,\chi_{j},\qquad 1\leq j\leq\mathbf{N},\quad 0\leq l\leq 2\mathbf{N}.\end{cases}$ (5.27)
From (5.27) we see that we have
$m_{l}(k_{j}^{\ast})=m_{l}(k_{j})^{\ast},\qquad 1\leq j\leq\mathbf{N}.$
We emphasize that we suppress the $t$-dependence in our notation for $\mathbf{M}(x),$ $\mathbf{A}(x),$ and $m_{l}(k_{j}).$ It turns out that the potential $Q$ is determined by $A_{2\mathbf{N}-1}(x)$ alone and the potential $P$ is determined by $A_{2\mathbf{N}-1}(x)$ and $A_{2\mathbf{N}-2}(x)$ only. The remaining entries $A_{j}(x)$ for $0\leq j\leq 2\mathbf{N}-3,$ along with $A_{2\mathbf{N}-1}(x)$ and $A_{2\mathbf{N}-2}(x),$ are used to construct the basic solutions $f(k,x),$ $g(k,x),$ $m(k,x),$ and $n(k,x)$ to (2.8) given in (5.11)–(5.14). It is understood that when $\mathbf{N}=1,$ we only have $A_{1}(x)$ and $A_{0}(x)$ and those two quantities determine both the potentials and the basic solutions.
In order to obtain $A_{j}(x)$ for $0\leq j\leq 2\mathbf{N}-1,$ we solve the linear system in (5.24) and obtain $\mathbf{A}(x)$ as
$\mathbf{A}(x)=-\mathbf{M}(x)^{-1}\mathbf{B}(x).$ (5.28)
From (5.28) we recover $A_{2\mathbf{N}-1}(x)$ and $A_{2\mathbf{N}-2}(x)$ as
$A_{2\mathbf{N}-1}(x)=-\begin{bmatrix}1&0&0&\cdots&0\end{bmatrix}\mathbf{M}(x)^{-1}\mathbf{B}(x),$ (5.29)
$A_{2\mathbf{N}-2}(x)=-\begin{bmatrix}0&1&0&\cdots&0\end{bmatrix}\mathbf{M}(x)^{-1}\mathbf{B}(x),$ (5.30)
where the row vector in (5.29) has $1$ in the first entry and the zeros in each of the remaining $2\mathbf{N}-1$ entries. Similarly, the row vector in (5.30) has $1$ in the second entry and the zeros in all remaining entries. By using (15) on p.12 of [9], from (5.29) and (5.30) we obtain $A_{2\mathbf{N}-1}(x)$ and $A_{2\mathbf{N}-2}(x)$ as the ratio of two determinants as
$A_{2\mathbf{N}-1(x)}=\displaystyle\frac{\det\begin{bmatrix}\begin{array}[]{c | c c c c c}0&1&0&0&\cdots&0\\ \hline\cr\mathbf{B}(x)&&&&\mathbf{M}(x)&\end{array}\end{bmatrix}}{\det\begin{bmatrix}\mathbf{M}(x)\end{bmatrix}}$,
$A_{2\mathbf{N}-2(x)}=\displaystyle\frac{\det\begin{bmatrix}\begin{array}[]{c | c c c c c}0&0&1&0&\cdots&0\\ \hline\cr\mathbf{B}(x)&&&&\mathbf{M}(x)&\end{array}\end{bmatrix}}{\det\begin{bmatrix}\mathbf{M}(x)\end{bmatrix}}$.
Alternatively, by using Cramer’s rule, from (5.24) we get
$A_{2\mathbf{N}-1}(x)=-\displaystyle\frac{\det[\mathbf{M}_{1}(x)]}{\det[\mathbf{M}(x)]},\qquad\mathbf{N}\geq 1,$ (5.31)
$A_{2\mathbf{N}-2}(x)=-\displaystyle\frac{\det[\mathbf{M}_{2}(x)]}{\det[\mathbf{M}(x)]},\qquad\mathbf{N}\geq 1.$ (5.32)
Here, $\mathbf{M}_{1}(x)$ denotes the $2\mathbf{N}\times 2\mathbf{N}$ matrix obtained by replacing the first column of the matrix $\mathbf{M}(x)$ with the column vector $\mathbf{B}(x),$ and $\mathbf{M}_{2}(x)$ denotes the $2\mathbf{N}\times 2\mathbf{N}$ matrix obtained by replacing the second column of $\mathbf{M}(x)$ with $\mathbf{B}(x).$
In order to solve the inverse scattering problem for (2.8) in the reflectionless case, we recall that we use the input data set $\{\eta_{j},\gamma_{j}\}_{j=1}^{\mathbf{N}},$ which allows a unique determination of the potentials $Q$ and $P.$ From (5.27), we see that all entries in the matrix $\mathbf{M}(x)$ and all components in the column vector $\mathbf{B}(x)$ are uniquely determined by the elements of the set $\{\eta_{j},\gamma_{j}\}_{j=1}^{\mathbf{N}}.$ Hence, as shown in (5.28) we uniquely determine $\mathbf{A}(x)$ by using the input data set $\{\eta_{j},\gamma_{j}\}_{j=1}^{\mathbf{N}}.$ In particular, we determine $A_{2\mathbf{N}-1}(x)$ and $A_{2\mathbf{N}-2}(x)$ uniquely in terms of the elements in the input set $\{\eta_{j},\gamma_{j}\}_{j=1}^{\mathbf{N}}.$
Having determined $A_{2\mathbf{N}-1}(x)$ and $A_{2\mathbf{N}-2}(x)$ uniquely by using the input data set $\{\eta_{j},\gamma_{j}\}_{j=1}^{\mathbf{N}},$ we proceed to determine the potentials $Q$ and $P$ appearing in (2.8). Using (5.2), (5.11), and (5.13) we obtain the asymptotics of $f(k,x)$ as $k\to\infty\text{\rm{ in }}\overline{\Omega_{1}}$ as
$f(k,x)=e^{kx}\left[1+\displaystyle\frac{A_{2\mathbf{N}-1}(x)-\Sigma_{\mathbf{N}}}{k}+\displaystyle\frac{A_{2\mathbf{N}-2}(x)-\Sigma_{\mathbf{N}}A_{2\mathbf{N}-1}(x)+\Pi_{\mathbf{N}}}{k^{2}}+O\left(\displaystyle\frac{1}{k^{3}}\right)\right],$ (5.33)
and the asymptotics of $g(k,x)$ as $k\to\infty\text{\rm{ in }}\overline{\Omega_{3}}$ as
$g(k,x)=e^{kx}\left[1+\displaystyle\frac{A_{2\mathbf{N}-1}(x)+\Sigma_{\mathbf{N}}}{k}+\displaystyle\frac{A_{2\mathbf{N}-2}(x)+\Sigma_{\mathbf{N}}A_{2\mathbf{N}-1}(x)+\Pi_{\mathbf{N}}}{k^{2}}+O\left(\displaystyle\frac{1}{k^{3}}\right)\right].$ (5.34)
We note that $\Pi_{\mathbf{N}}$ appearing in (5.33) and (5.34) is the constant defined as
$\begin{split}\Pi_{\mathbf{N}}:=&k_{1}\left(k_{1}+k_{1}^{\ast}+\cdots+k_{\mathbf{N}}+k_{\mathbf{N}}^{\ast}\right)+k_{1}^{\ast}\left(k_{1}^{\ast}+k_{2}+\cdots+k_{\mathbf{N}}+k_{\mathbf{N}}^{\ast}\right)\\ &+k_{2}\left(k_{2}+k_{2}^{\ast}+\cdots+k_{\mathbf{N}}+k_{\mathbf{N}}^{\ast}\right)+\cdots+k_{\mathbf{N}}\left(k_{\mathbf{N}}+k_{\mathbf{N}}^{\ast}\right)+k_{\mathbf{N}}^{\ast}\left(k_{\mathbf{N}}^{\ast}\right),\end{split}$
and we recall that $\Sigma_{\mathbf{N}}$ is the constant defined in (5.3). Comparing (3.6) with (5.33) and comparing (3.7) with (5.34), we obtain
$u_{1}(x)=A_{2\mathbf{N}-1}(x)-\Sigma_{\mathbf{N}},$ (5.35)
$v_{1}(x)=A_{2\mathbf{N}-1}(x)+\Sigma_{\mathbf{N}},$ (5.36)
$u_{2}(x)=A_{2\mathbf{N}-2}(x)-\Sigma_{\mathbf{N}}A_{2\mathbf{N}-1}(x)+\Pi_{\mathbf{N}},$ (5.37)
$v_{2}(x)=A_{2\mathbf{N}-2}(x)+\Sigma_{\mathbf{N}}A_{2\mathbf{N}-1}(x)+\Pi_{\mathbf{N}}.$ (5.38)
With the help of (3.14)–(3.17) and (5.35)–(5.38), we get
$Q(x)=-3\,\displaystyle\frac{dA_{2\mathbf{N}-1}(x)}{dx},$ (5.39)
$P(x)=3\left(A_{2\mathbf{N}-1}(x)\,\displaystyle\frac{dA_{2\mathbf{N}-1}(x)}{dx}-\displaystyle\frac{d^{2}A_{2\mathbf{N}-1}(x)}{dx^{2}}-\displaystyle\frac{dA_{2\mathbf{N}-2}(x)}{dx}\right).$ (5.40)
Since $A_{2\mathbf{N}-1}(x)$ and $A_{2\mathbf{N}-2}(x)$ are uniquely determined by the input data set $\{\eta_{j},\gamma_{j}\}_{j=1}^{\mathbf{N}},$ we see from (5.39) and (5.40) that the same input data set uniquely determines the potentials $Q$ and $P$ appearing in (2.8). With the help of the first lines of (3.4) and (3.5), from (5.33) and (5.34) we obtain the spacial asymptotics given by
$\lim_{x\to\pm\infty}A_{2\mathbf{N}-1}(x)=\pm\Sigma_{\mathbf{N}},$ (5.41)
$\lim_{x\to\pm\infty}A_{2\mathbf{N}-2}(x)=\Sigma_{\mathbf{N}}^{2}-\Pi_{\mathbf{N}}.$ (5.42)
We remark that the determinants of the matrices $\mathbf{M}(x),$ $\mathbf{M}_{1}(x),$ and $\mathbf{M}_{2}(x)$ appearing in (5.31) and (5.32) have various useful properties. Next, we elaborate on some of those properties. For example, the determinant of $\mathbf{M}_{1}(x)$ is determined in terms of the determinant of $\mathbf{M}(x)$ as
$\det[\mathbf{M}_{1}(x)]=\Sigma_{\mathbf{N}}\det[\mathbf{M}(x)]+\displaystyle\frac{d}{dx}\det[\mathbf{M}(x)],$ (5.43)
which enables us to write (5.31) in the equivalent form as
$A_{2\mathbf{N}-1}(x)=-\Sigma_{\mathbf{N}}-\displaystyle\frac{1}{\det[\mathbf{M}(x)]}\displaystyle\frac{d}{dx}\det[\mathbf{M}(x)].$ (5.44)
Consequently, using (5.44) on the right-hand side of (5.39), we express $Q$ in terms of the determinant of $\mathbf{M}(x)$ as
$Q(x)=3\,\displaystyle\frac{d}{dx}\left(\displaystyle\frac{1}{\det[\mathbf{M}(x)]}\,\displaystyle\frac{d\det[\mathbf{M}(x)]}{dx}\right).$ (5.45)
As seen from (5.22) and (5.25)–(5.27), the appearance of the parameters $x$ and $t$ is restricted through the quantities $\chi_{j}$ appearing in (5.20). In fact, from (5.25)–(5.27) we observe that each of the determinants of the matrices $\mathbf{M}(x),$ $\mathbf{M}_{1}(x),$ and $\mathbf{M}_{2}(x)$ is a polynomial of degree $2\mathbf{N}$ in the $\mathbf{N}$ real-valued variables $\chi_{1},\dots,\chi_{\mathbf{N}}.$ Because of the specific appearance of $k_{j}$ and $k_{j}^{\ast}$ in (5.2) for $1\leq j\leq\mathbf{N},$ the variables $\chi_{1},\dots,\chi_{\mathbf{N}}$ appear on those three determinants in a symmetrical way. Each variable $\chi_{j}$ appears in the polynomial at most quadratically. The aforementioned symmetry allows us to determine the coefficients in those polynomials explicitly in terms of the parameters in the set $\{k_{j},k_{j}^{\ast},\gamma_{j},\gamma_{j}^{\ast}\}_{j=1}^{\mathbf{N}}.$ For example, if we determine the coefficient of the term $\chi_{1}$ in the polynomial, then we also get the coefficient of each $\chi_{j}$ as a result of the aforementioned symmetry. By setting $\chi_{j}=0$ for $2\leq j\leq\mathbf{N},$ from (5.25) we get
$\det[\mathbf{M}(x)]=\det\begin{bmatrix}k_{1}^{2\mathbf{N}-1}+(zk_{1})^{2\mathbf{N}-1}\gamma_{1}\,\chi_{1}&\cdots&k_{1}+zk_{1}\gamma_{1}\,\chi_{1}&1+\gamma_{1}\,\chi_{1}\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr(k_{1}^{\ast})^{2\mathbf{N}-1}+(z^{2}k_{1}^{\ast})^{2\mathbf{N}-1}\gamma_{1}^{\ast}\,\chi_{1}&\cdots&k_{1}^{\ast}+z^{2}k_{1}^{\ast}\gamma_{1}^{\ast}\chi_{1}&1+\gamma_{1}^{\ast}\chi_{1}\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr\vdots&\ddots&\vdots&\vdots\\ k_{\mathbf{N}}^{2\mathbf{N}-1}&\cdots&k_{\mathbf{N}}&1\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr(k_{\mathbf{N}}^{\ast})^{2\mathbf{N}-1}&\cdots&k_{\mathbf{N}}^{\ast}&1\end{bmatrix}$, (5.46)
which can be written as
$\det[\mathbf{M}(x)]=\alpha_{0}+\alpha_{11}\,\chi_{1}+\alpha_{12}\,\chi_{1}^{2},$ (5.47)
where the coefficients $\alpha_{0},$ $\alpha_{11},$ and $\alpha_{12}$ are explicitly expressed in terms of the parameters in the set $\{k_{j},k_{j}^{\ast},\gamma_{j},\gamma_{j}^{\ast}\}_{j=1}^{\mathbf{N}}.$ The coefficient $\alpha_{0}$ is obtained by letting $\chi_{1}=0$ in (5.46), from which we obtain $\alpha_{0}$ as
$\alpha_{0}=\det\begin{bmatrix}k_{1}^{2\mathbf{N}-1}&k_{1}^{2\mathbf{N}-2}&\cdots&k_{1}&1\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr(k_{1}^{\ast})^{2\mathbf{N}-1}&(k_{1}^{\ast})^{2\mathbf{N}-2}&\cdots&k_{1}^{\ast}&1\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr\vdots&\vdots&\ddots&\vdots&\vdots\\ k_{\mathbf{N}}^{2\mathbf{N}-1}&k_{\mathbf{N}}^{2\mathbf{N}-2}&\cdots&k_{\mathbf{N}}&1\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr(k_{\mathbf{N}}^{\ast})^{2\mathbf{N}-1}&(k_{\mathbf{N}}^{\ast})^{2\mathbf{N}-2}&\cdots&k_{\mathbf{N}}^{\ast}&1\end{bmatrix}$. (5.48)
The right-hand side in (5.48) is related to a Vandermonde determinant, and we write (5.48) as
$\alpha_{0}=\mathcal{V}(k_{1},k_{1}^{\ast},k_{2},k_{2}^{\ast},\cdots,k_{\mathbf{N}},k_{\mathbf{N}}^{\ast}),$ (5.49)
with the Vandermonde coefficient on the right-hand side of (5.49) is defined as
$\mathcal{V}(a_{1},a_{2},a_{3},a_{4},\cdots,a_{2p-1},a_{2p}):=\prod_{1\leq j<l\leq 2p}(a_{j}-a_{l}).$ (5.50)
We write (5.49) as
$\alpha_{0}=\mathcal{U}_{00},$ (5.51)
and we recover the coefficient $\alpha_{11}$ appearing in (5.47) from (5.46) as
$\alpha_{11}=\gamma_{1}\,\mathcal{U}_{10}+\gamma_{1}^{\ast}\,\mathcal{U}_{01}.$ (5.52)
We also obtain the coefficient $\alpha_{12}$ in (5.47) as
$\alpha_{12}=\gamma_{1}\gamma_{1}^{\ast}\,\mathcal{U}_{11},$ (5.53)
where we have defined
$\begin{cases}\mathcal{U}_{00}:=\mathcal{V}(k_{1},k_{1}^{\ast},k_{2},k_{2}^{\ast},\cdots,k_{\mathbf{N}},k_{\mathbf{N}}^{\ast}),\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr\mathcal{U}_{10}:=\mathcal{V}(zk_{1},k_{1}^{\ast},k_{2},k_{2}^{\ast},\cdots,k_{\mathbf{N}},k_{\mathbf{N}}^{\ast}),\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr\mathcal{U}_{01}:=\mathcal{V}(k_{1},z^{2}k_{1}^{\ast},k_{2},k_{2}^{\ast},\cdots,k_{\mathbf{N}},k_{\mathbf{N}}^{\ast}),\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr\mathcal{U}_{11}:=\mathcal{V}(zk_{1},z^{2}k_{1}^{\ast},k_{2},k_{2}^{\ast},\cdots,k_{\mathbf{N}},k_{\mathbf{N}}^{\ast}).\end{cases}$ (5.54)
In a similar manner, by setting $\chi_{j}=0$ for $2\leq j\leq\mathbf{N}$ in the expression for $\det[\mathbf{M}_{2}(x)],$ we get
$\det[\mathbf{M}_{2}(x)]=\beta_{0}+\beta_{11}\,\chi_{1}+\beta_{12}\,\chi_{1}^{2}.$ (5.55)
During the evaluation of $\det[\mathbf{M}_{2}(x)],$ we can factor out certain common terms from the rows of the matrix $\mathbf{M}_{2}(x)$ and then evaluate the determinant of the resulting simplified matrix as a Vandermonde determinant. Hence, in order to express the coefficients $\beta_{0},$ $\beta_{11},$ $\beta_{12}$ in (5.55) explicitly in terms of the parameters in the set $\{k_{j},k_{j}^{\ast},\gamma_{j},\gamma_{j}^{\ast}\}_{j=1}^{\mathbf{N}},$ it is convenient to introduce the modified Vandermonde coefficient $\tilde{\mathcal{V}}(a_{1},a_{2},a_{3},a_{4},\cdots,a_{2p-1},a_{2p})$ as
$\tilde{\mathcal{V}}(a_{1},a_{2},a_{3},a_{4},\cdots,a_{2p-1},a_{2p}):=\left(\sum_{1\leq j<l\leq 2p}a_{j}\,a_{l}\right)\mathcal{V}(a_{1},a_{2},a_{3},a_{4},\cdots,a_{2p-1},a_{2p}).$
With the help of the modified Vandermonde coefficient, we get
$\beta_{0}=-\tilde{\mathcal{V}}(k_{1},k_{1}^{\ast},k_{2},k_{2}^{\ast},\cdots,k_{\mathbf{N}},k_{\mathbf{N}}^{\ast}),$
$\beta_{11}=-\gamma_{1}\,\tilde{\mathcal{V}}(zk_{1},k_{1}^{\ast},k_{2},k_{2}^{\ast},\cdots,k_{\mathbf{N}},k_{\mathbf{N}}^{\ast})-\gamma_{1}^{\ast}\,\tilde{\mathcal{V}}(k_{1},z^{2}k_{1}^{\ast},k_{2},k_{2}^{\ast},\cdots,k_{\mathbf{N}},k_{\mathbf{N}}^{\ast}),$
$\beta_{12}=-\gamma_{1}\,\gamma_{1}^{\ast}\,\tilde{\mathcal{V}}(zk_{1},z^{2}k_{1}^{\ast},k_{2},k_{2}^{\ast},\cdots,k_{\mathbf{N}},k_{\mathbf{N}}^{\ast}).$
As mentioned already, the determinant of $\mathbf{M}(x)$ is a polynomial of degree $2\mathbf{N}$ in the $\mathbf{N}$ real-valued variables $\chi_{1},\dots,\chi_{\mathbf{N}}.$ The coefficient of the highest-degree term $\chi_{1}^{2}\chi_{2}^{2}\cdots\chi_{\mathbf{N}}^{2}$ in $\det[\mathbf{M}(x)]$ is obtained in a manner similar to the evaluation of $\alpha_{12}$ given in (5.53), and that coefficient is given by
$\left(\prod_{j=1}^{\mathbf{N}}\gamma_{j}\,\gamma_{j}^{\ast}\right)\mathcal{V}(zk_{1},z^{2}k_{1}^{\ast},zk_{2},z^{2}k_{2}^{\ast},\cdots,zk_{\mathbf{N}},z^{2}k_{\mathbf{N}}^{\ast}).$ (5.56)
Since $\det[\mathbf{M}_{1}(x)]$ is also a polynomial of degree $2\mathbf{N}$ in the $\mathbf{N}$ real-valued variables $\chi_{1},\dots,\chi_{\mathbf{N}},$ as in (5.47) and (5.55), by setting $\chi_{j}=0$ for $2\leq j\leq\mathbf{N}$ in the expression for $\det[\mathbf{M}_{1}(x)],$ we obtain
$\det[\mathbf{M}_{1}(x)]=\epsilon_{0}+\epsilon_{11}\,\chi_{1}+\epsilon_{12}\,\chi_{1}^{2}.$ (5.57)
With the help of (5.20), (5.43), and (5.47), we evaluate the coefficients in (5.57) as
$\epsilon_{0}=\Sigma_{\mathbf{N}}\alpha_{0},\quad\epsilon_{11}=[\Sigma_{\mathbf{N}}+(z-1)k_{1}]\alpha_{11},\quad\epsilon_{12}=[\Sigma_{\mathbf{N}}+2(z-1)k_{1}]\alpha_{12},$
where we recall that $\Sigma_{\mathbf{N}},$ $\alpha_{11},$ and $\alpha_{12}$ are the quantities in (5.3), (5.52), and (5.53), respectively.
As already indicated, the symmetrical appearance of the quantities $k_{j}$ in (5.2) for $1\leq j\leq N$ allows us to determine the coefficients of $\chi_{j}$ and $\chi_{j}^{2}$ in the expressions for $\det[\mathbf{M}(x)],$ $\det[\mathbf{M}_{1}(x)],$ and $\det[\mathbf{M}_{2}(x)].$ For example, with $\chi_{j}\neq 0$ if we choose the remaining quantities $\chi_{l}$ as zero for $l\neq j,$ then we get the analog of (5.47) as
$\det[\mathbf{M}(x)]=\alpha_{0}+\alpha_{jj}\,\chi_{j}+\alpha_{j(j+1)}\,\chi_{j}^{2},$ (5.58)
where $\alpha_{jj}$ and $\alpha_{j(j+1)}$ are the analogs of $\alpha_{11}$ and $\alpha_{12}$ in (5.52) and (5.53), respectively, and they are expressed as
$\alpha_{jj}=\gamma_{j}\,\mathcal{U}_{j0}+\gamma_{j}^{\ast}\,\mathcal{U}_{0j},\quad\alpha_{j(j+1)}=\gamma_{j}\,\gamma_{j}^{\ast}\,\mathcal{U}_{jj},\qquad 1\leq j\leq\mathbf{N}.$ (5.59)
We remark that $\mathcal{U}_{j0}$ for $2\leq j\leq\mathbf{N}$ is defined in a similar way $\mathcal{U}_{10}$ appearing in the second line of (5.54) is defined, i.e. by replacing $k_{j}$ appearing on the right-hand side of (5.49) by $zk_{j}.$ Similarly, $\mathcal{U}_{0j}$ is defined in a similar way $\mathcal{U}_{01}$ appearing in the third line of (5.54) is defined, i.e. by replacing $k_{j}^{\ast}$ appearing on the right-hand side of (5.49) by $z^{2}k_{j}^{\ast}.$ Likewise, $\mathcal{U}_{jj}$ is defined in a similar way $\mathcal{U}_{11}$ appearing in the fourth line of (5.54) is defined, i.e. by replacing $k_{j}$ and $k_{j}^{\ast}$ appearing on the right-hand side of (5.49) by $zk_{j}$ and $z^{2}k_{j}^{\ast},$ respectively.
6 Soliton solutions to the Sawada–Kotera equation
In Section 5 we have constructed complex-valued solution pairs $(Q,P)$ to the integrable system (2.6). This is done by solving the inverse scattering problem for (2.8) in the reflectionless case using the bound-state poles of the left transmission coefficient $T_{\text{\rm{l}}}(k)$ given in (5.1) and the time-evolved modified bound-state dependency constants. By using the input data set $\{k_{j},\gamma_{j}\}_{j=1}^{\mathbf{N}},$ with $k_{j}$ being the quantity in (5.4) and $\gamma_{j}$ being the complex-valued modified bound-state dependency constant in (5.15), we have constructed the potential pair $(Q,P)$ as in (5.39) and (5.40). Since we have used the time evolution for the dependency constants as specified in (4.5) and (5.18), the resulting time-evolved potential pair $(Q,P)$ solves the integrable coupled system (2.6).
In this section, by imposing the appropriate restriction on the solution pair $(Q,P)$ appearing in (5.39) and (5.40), we construct the $\mathbf{N}$-soliton solution to the Sawada–Kotera equation (2.1). It turns out that relevant restriction is accomplished by specifying each of the ratios $s_{j}/r_{j}$ for $1\leq j\leq\mathbf{N}$ in the input data set $\{k_{j},\gamma_{j}\}_{j=1}^{\mathbf{N}}$ appropriately in terms of the elements in the set $\{k_{j}\}_{j=1}^{\mathbf{N}}$ or equivalently in the set $\{\eta_{j}\}_{j=1}^{\mathbf{N}}.$ Here, we use $r_{j}$ and $s_{j}$ to denote the real and imaginary parts of $\gamma_{j},$ respectively. Thus, we have
$\gamma_{j}=r_{j}+is_{j},\qquad 1\leq j\leq\mathbf{N}.$ (6.1)
Because of the symmetrical appearance of the quantities $k_{j}$ in (5.2), for the determination of $s_{j}/r_{j}$ for $1\leq j\leq\mathbf{N},$ it is sufficient to determine the explicit expression for $s_{1}/r_{1}$ and use the appropriate symmetries to determine the remaining ratios $s_{j}/r_{j}.$ The restrictions on the ratios $s_{j}/r_{j}$ for $1\leq j\leq\mathbf{N}$ assure that $Q$ is real valued and we have either $P\equiv 0$ or $P=Q_{x}.$ The resulting real-valued expression for $Q$ contains the $2\mathbf{N}$ real-valued parameters $\eta_{j}$ and $r_{j}$ for $1\leq j\leq\mathbf{N},$ where we recall that $\eta_{j}$ is related to $k_{j}$ as in (5.4).
In the following steps, we show how the real-valued $\mathbf{N}$-soliton solution $Q$ to (2.1) expressed in terms of the parameters in the set $\{\eta_{j},r_{j}\}_{j=1}^{\mathbf{N}}$ is obtained from the complex-valued solution pair $(Q,P)$ in (5.39) and (5.40) expressed in terms of the parameters in the set $\{\eta_{j},r_{j},s_{j}\}_{j=1}^{\mathbf{N}}.$
  1. (a)
    When $P\equiv 0,$ by integrating both sides of (5.40) and using (5.41) and (5.42), we write (5.40) in the equivalent form as
    $\frac{1}{2}\left[A_{2\mathbf{N}-1}(x)\right]^{2}-A^{\prime}_{2\mathbf{N}-1}(x)-A_{2\mathbf{N}-2}(x)=\Pi_{\mathbf{N}}-\displaystyle\frac{1}{2}\,\Sigma_{\mathbf{N}}^{2},$ (6.2)
    where we recall that we use a prime to denote the $x$-derivative. Using (5.32) and (5.44) in (6.2), after some simplifications, we obtain the equivalent expression given by
    $\begin{split}\left(\Pi_{\mathbf{N}}-\Sigma_{\mathbf{N}}^{2}\right)\det[\mathbf{M}(x)]+\Sigma_{\mathbf{N}}\,\displaystyle\frac{d\det[\mathbf{M}(x)]}{dx}+&\displaystyle\frac{d^{2}\det[\mathbf{M}(x)]}{dx^{2}}+\det[\mathbf{M}_{2}(x)]\\ &=\displaystyle\frac{1}{2\,\det[\mathbf{M}(x)]}\left[\displaystyle\frac{d\det[\mathbf{M}(x)]}{dx}\right]^{2}.\end{split}$ (6.3)
  2. (b)
    On the other hand, if $P=Q_{x}$ then (5.39) and (5.40) yield
    $A_{2\mathbf{N}-1}(x)\,A^{\prime}_{2\mathbf{N}-1}(x)-A^{\prime}_{2\mathbf{N}-2}(x)=0.$ (6.4)
    Integrating both sides of (6.4) and using (5.39), (5.41), and (5.42), we write (6.4) in the equivalent form as
    $\frac{1}{2}\left[A_{2\mathbf{N}-1}(x)\right]^{2}-A_{2\mathbf{N}-2}(x)=\Pi_{\mathbf{N}}-\displaystyle\frac{1}{2}\,\Sigma_{\mathbf{N}}^{2}.$ (6.5)
    Using (5.32) and (5.44) in (6.5), after some simplifications, we write (6.5) in the equivalent form given by
    $\begin{split}\left(\Pi_{\mathbf{N}}-\Sigma_{\mathbf{N}}^{2}\right)\det[\mathbf{M}(x)]+\Sigma_{\mathbf{N}}\,\displaystyle\frac{d\det[\mathbf{M}(x)]}{dx}+&\displaystyle\frac{d^{2}\det[\mathbf{M}(x)]}{dx^{2}}+\det[\mathbf{M}_{2}(x)]\\ &=-\displaystyle\frac{1}{2\,\det[\mathbf{M}(x)]}\left[\displaystyle\frac{d\det[\mathbf{M}(x)]}{dx}\right]^{2}.\end{split}$ (6.6)
  3. (c)
    From Section 5, we know that each of $\det[\mathbf{M}(x)]$ and $\det[\mathbf{M}_{2}(x)]$ is a polynomial of degree $2\mathbf{N}$ in the $\mathbf{N}$ real-valued variables $\chi_{1},\dots,\chi_{\mathbf{N}}.$ Consequently, each of $d\det[\mathbf{M}(x)]/dx$ and $d^{2}\det[\mathbf{M}(x)]/dx^{2}$ is also a polynomial of degree $2\mathbf{N}$ in $\chi_{1},\dots,\chi_{\mathbf{N}}.$ Thus, the left-hand sides of (6.3) and (6.6) are some polynomials of degree $2\mathbf{N}$ in $\chi_{1},\dots,\chi_{\mathbf{N}}.$ We remark that the left-hand sides of (6.3) and (6.6) coincide while their right-hand sides differ by a minus sign. Thus, those right-hand sides must also be polynomials of degree $2\mathbf{N}$ in $\chi_{1},\dots,\chi_{\mathbf{N}}.$
  4. (d)
    In order to satisfy the restrictions that the right-hand sides of (6.3) and (6.6) must each be a polynomial of degree $2\mathbf{N}$ in $\chi_{1},\dots,\chi_{\mathbf{N}},$ we let
    $\det[\mathbf{M}(x)]=\alpha_{0}\left[\Delta(x)\right]^{2},$ (6.7)
    where $\alpha_{0}$ is the constant appearing in (5.47)–(5.49) and $\Delta(x)$ is a polynomial of degree $\mathbf{N}$ in $\chi_{1},\dots,\chi_{\mathbf{N}}.$ The use of $\alpha_{0}$ in (6.7) ensures that $\Delta(x)$ is normalized in the sense that $\Delta(x)=1$ when we choose $\chi_{j}=0$ for $1\leq j\leq\mathbf{N}.$ Comparing (5.47) with (6.7), we see that, if we set $\chi_{j}=0$ for $2\leq j\leq\mathbf{N},$ then $\Delta(x)$ must satisfy
    $\Delta(x)]^{2}=1+\displaystyle\frac{\alpha_{11}}{\alpha_{0}}\,\chi_{1}+\displaystyle\frac{\alpha_{12}}{\alpha_{0}}\,\chi_{1}^{2},$ (6.8)
    and hence the coefficients $\alpha_{0},$ $\alpha_{11},$ and $\alpha_{12}$ appearing in (5.47) must be related to each other as
    $\frac{\alpha_{12}}{\alpha_{0}}=\left(\displaystyle\frac{\alpha_{11}}{2\,\alpha_{0}}\right)^{2},$ (6.9)
    or equivalently as
    $4\,\alpha_{0}\,\alpha_{12}=\left[\alpha_{11}\right]^{2}.$ (6.10)
  5. (e)
    From (5.51)–(5.54) we observe that the constraint (6.10) can be expressed in terms of the Vandermonde coefficients appearing in (5.54) as
    $4\,\mathcal{U}_{00}\,\mathcal{U}_{11}\,\gamma_{1}\,\gamma_{1}^{\ast}=\left(\mathcal{U}_{10}\,\gamma_{1}+\mathcal{U}_{01}\,\gamma_{1}^{\ast}\right)^{2}.$ (6.11)
    Using (6.1) in (6.11), we write (6.11) in terms of $r_{1}$ and $s_{1}$ as
    $4\,\mathcal{U}_{00}\,\mathcal{U}_{11}\left(r_{1}^{2}+s_{1}^{2}\right)=\left[\mathcal{U}_{10}\,(r_{1}+is_{1})+\mathcal{U}_{01}\,(r_{1}-is_{1})\right]^{2}.$ (6.12)
    We divide both sides of (6.12) by $r_{1}^{2}$ and write it in the equivalent form expressed as a quadratic equation in $s_{1}/r_{1}$ as
    $\begin{split}\left[4\,\mathcal{U}_{00}\,\mathcal{U}_{11}+(\mathcal{U}_{10}-\mathcal{U}_{01})^{2}\right]\left(\displaystyle\frac{s_{1}}{r_{1}}\right)^{2}-&2i\left(\mathcal{U}_{10}^{2}-\mathcal{U}_{01}^{2}\right)\left(\displaystyle\frac{s_{1}}{r_{1}}\right)\\ &+\left[4\,\mathcal{U}_{00}\,\mathcal{U}_{11}-(\mathcal{U}_{10}+\mathcal{U}_{01})^{2}\right]=0.\end{split}$ (6.13)
    The two solutions to the quadratic equation (6.13) are given by
    $\frac{s_{1}}{r_{1}}=\displaystyle\frac{i\,(\mathcal{U}_{10}^{2}-\mathcal{U}_{01}^{2})\pm 4\sqrt{\mathcal{U}_{00}\,\mathcal{U}_{11}\,(\mathcal{U}_{10}\,\mathcal{U}_{01}-\mathcal{U}_{00}\,\mathcal{U}_{11})}}{4\,\mathcal{U}_{00}\,\mathcal{U}_{11}+(\mathcal{U}_{10}-\mathcal{U}_{01})^{2}}.$ (6.14)
    From (5.50) and (5.55) it follows that the denominator on the right-hand side is real and nonzero, the quantity inside the square root in the numerator is positive, and $(\mathcal{U}_{10}^{2}-\mathcal{U}_{01}^{2})$ is purely imaginary. Hence, the quadratic equation (6.13) has two distinct real roots specified in (6.14). In fact, one of them is related to the constraint regarding the right-hand side of (6.3) and the other is related to the constraint regarding the right-hand side of (6.6).
  6. (f)
    Exploiting the symmetries existing in (5.47) and (5.58), with the help of (5.59), from (6.14) we obtain the ratios $s_{j}/r_{j}$ as
    $\frac{s_{j}}{r_{j}}=\displaystyle\frac{i\,(\mathcal{U}_{j0}^{2}-\mathcal{U}_{0j}^{2})\pm 4\sqrt{\mathcal{U}_{00}\,\mathcal{U}_{jj}\,(\mathcal{U}_{j0}\,\mathcal{U}_{0j}-\mathcal{U}_{00}\,\mathcal{U}_{jj})}}{4\,\mathcal{U}_{00}\,\mathcal{U}_{jj}+(\mathcal{U}_{j0}-\mathcal{U}_{0j})^{2}},\qquad 1\leq j\leq\mathbf{N}.$ (6.15)
    As indicated in (e), the two root values appearing on the right-hand side of (6.15) are real and distinct.
  7. (g)
    We already know that each of $\mathcal{U}_{00},$ $\mathcal{U}_{j0},$ $\mathcal{U}_{0j},$ $\mathcal{U}_{jj}$ is determined in terms of the elements in the set $\{k_{j}\}_{j=1}^{\mathbf{N}}$ or equivalently in $\{\eta_{j}\}_{j=1}^{\mathbf{N}}$ due to (5.4). From (6.15) we observe that each ratio $s_{j}/r_{j}$ is uniquely expressed in terms of the elements in $\{\eta_{j}\}_{j=1}^{\mathbf{N}}$ in each of the cases $P\equiv 0$ and $P=Q_{x}.$ Thus, each of the quantities $(1+is_{j}/r_{j})$ is a constant uniquely expressed in terms of the elements in $\{\eta_{j}\}_{j=1}^{\mathbf{N}}.$ Let us write (6.1) as
    $\gamma_{j}=r_{j}\left(1+i\,\displaystyle\frac{s_{j}}{r_{j}}\right),\qquad 1\leq j\leq\mathbf{N}.$ (6.16)
    Using (6.15) on the right-hand side of (6.16), we express each $\gamma_{j}$ as a product of $r_{j}$ and a constant determined by $\{\eta_{j}\}_{j=1}^{\mathbf{N}}.$ By using the right-hand side of (6.16) on the right-hand side of (5.27), we express all entries of the matrix $\mathbf{M}(x)$ appearing on the right-hand side of (5.25) in terms of the elements in $\{\eta_{j},r_{j}\}_{j=1}^{\mathbf{N}}.$ Using the resulting expression for $\mathbf{M}(x)$ on the right-hand side of (5.45), we obtain the solution $Q$ to (2.1) in terms of the elements in $\{\eta_{j}\}_{j=1}^{\mathbf{N}}.$
  8. (h)
    Alternatively, we obtain $Q$ by proceeding as follows. Using either of the two roots specified in (6.15), we uniquely determine the quantity $\Delta(x)$ in terms of the parameters in the set $\{k_{j},r_{j}\}_{j=1}^{\mathbf{N}}$ or equivalently in $\{\eta_{j},r_{j}\}_{j=1}^{\mathbf{N}}.$ Using (6.7) in (5.45) we recover the solution $Q$ to (2.1) as
    $Q(x)=6\,\displaystyle\frac{d}{dx}\left[\displaystyle\frac{\Delta^{\prime}(x)}{\Delta(x)}\right],$ (6.17)
    where we recall that the prime is used for the $x$-derivative and we suppress the $t$-dependence in our notation $\Delta(x).$
  9. (i)
    We recall that $\Delta(x)$ is a real-valued polynomial of degree $\mathbf{N}$ in the $\mathbf{N}$ real variables $\chi_{1},\dots,\chi_{\mathbf{N}}.$ The constant term not containing any of those $\mathbf{N}$ variables is equal to $1.$ In each term, a specific $\chi_{j}$ either does not appear or appears only to the first power. From (5.51), (5.52), (6.1), (6.7)–(6.10) it follows that the coefficient of $\chi_{1}$ is equal to $\alpha_{11}/(2\alpha_{0}),$ where we have
    $\frac{\alpha_{11}}{2\alpha_{0}}=\displaystyle\frac{r_{1}}{2\,\mathcal{U}_{00}}\left[(\mathcal{U}_{10}+\mathcal{U}_{01})+i\,(\mathcal{U}_{10}-\mathcal{U}_{01})\,\displaystyle\frac{s_{1}}{r_{1}}\right],$ (6.18)
    with the understanding that (6.14) is used for the expression $s_{1}/r_{1}$ appearing on the right-hand side of (6.18). In a similar way, with the help of (5.58), (5.59), (6.1), (6.7)–(6.10) it follows that the coefficient of $\chi_{j}$ is equal to $\alpha_{jj}/(2\alpha_{0}),$ where we have
    $\frac{\alpha_{jj}}{2\alpha_{0}}=\displaystyle\frac{r_{j}}{2\,\mathcal{U}_{00}}\left[(\mathcal{U}_{j0}+\mathcal{U}_{0j})+i\,(\mathcal{U}_{j0}-\mathcal{U}_{0j})\,\displaystyle\frac{s_{j}}{r_{j}}\right],\qquad 1\leq j\leq\mathbf{N},$ (6.19)
    with the understanding that (6.15) is used for the expression $s_{j}/r_{j}$ appearing on the right-hand side of (6.19). The sign of the free real parameter $r_{j}$ is chosen so that the right-hand side of (6.19) is positive.
  10. (j)
    We introduce the quantities $y_{j}$ as
    $y_{j}:=\displaystyle\frac{\alpha_{jj}}{2\alpha_{0}}\,\chi_{j},\qquad 1\leq j\leq\mathbf{N},$ (6.20)
    where $\alpha_{jj}/(2\alpha_{0})$ is given by the right-hand side of (6.19) and we recall that $\chi_{j}$ is the quantity in (5.20). Using (5.4) in (5.20) we get
    $\chi_{j}=e^{\sqrt{3}\,\eta_{j}(x-9\eta_{j}^{4}t)},\qquad 1\leq j\leq\mathbf{N},$ (6.21)
    and hence $\chi_{j}$ remains positive for all real values of $x$ and $t.$ By choosing the sign of each $r_{j}$ appropriately so that the right-hand side of (6.19) is positive, we observe from (6.20) and (6.21) that the quantity $y_{j}$ defined in (6.20) remains positive for $1\leq j\leq\mathbf{N}$ for all $x$ and $t.$ We can then write $\Delta(x)$ in terms of $y_{j}$ as
    $\begin{split}\Delta(x)=&1+\sum_{j=1}^{\mathbf{N}}y_{j}+\sum_{1\leq j_{1}<j_{2}\leq\mathbf{N}}A_{j_{1}j_{2}}\,y_{j_{1}}\,y_{j_{2}}\\ &+\sum_{1\leq j_{1}<j_{2}<j_{3}\leq\mathbf{N}}A_{j_{1}j_{2}}\,A_{j_{1}j_{3}}\,A_{j_{2}j_{3}}\,y_{j_{1}}\,y_{j_{2}}\,y_{j_{3}}\\ &+\cdots+\left[A_{12}\,A_{13}\cdots\,A_{(\mathbf{N}-1)\mathbf{N}}\right]y_{1}\,y_{2}\cdots\,y_{\mathbf{N}},\end{split}$ (6.22)
    where the right-hand side is a polynomial in $y_{1},y_{2},\dots,y_{\mathbf{N}}$ with the last term containing the product $y_{1}y_{2}\cdots y_{\mathbf{N}}.$ We remark that (6.22) is the analog of (32) in [24]. With the help of (5.56) and (6.7), the double-indexed quantity $A_{jl}$ is determined by the elements in the set $\{k_{j}\}_{j=1}^{\mathbf{N}}$ and is given by
    $A_{jl}:=\displaystyle\frac{(k_{j}-k_{l})^{3}(k_{j}^{3}+k_{l}^{3})}{(k_{j}+k_{l})^{3}(k_{j}^{3}-k_{l}^{3})},\qquad 1\leq j<l\leq\mathbf{N}.$ (6.23)
    Using (5.4) in (6.23), we can express $A_{jl}$ in terms of the elements of $\{\eta_{j}\}_{j=1}^{\mathbf{N}}$ as
    $A_{jl}=\displaystyle\frac{(\eta_{j}-\eta_{l})^{3}(\eta_{j}^{3}+\eta_{l}^{3})}{(\eta_{j}+\eta_{l})^{3}(\eta_{j}^{3}-\eta_{l}^{3})},\qquad 1\leq j<l\leq\mathbf{N}.$ (6.24)
    We remark that (6.24) is the analog of (33) in [24]. Because the $\eta_{j}$-values are distinct, from (6.24) it follows that $A_{jl}$ is positive. Consequently, from (6.22) we observe that the quantity $\Delta(x)$ remains positive for all real values of $x$ and $t$ and this assures that the $\mathbf{N}$-soliton solution $Q$ to (2.1) appearing in (6.17) and containing the $2\mathbf{N}$ real-valued parameters $\eta_{j}$ and $r_{j}$ for $1\leq j\leq\mathbf{N}$ does not have any singularities.
7 Soliton behaviors
In Section 6 we have presented the construction of the $\mathbf{N}$-soliton solution to the Sawada–Kotera equation (2.1) with the input consisting of the $2\mathbf{N}$ real parameters in the set $\{\eta_{j},r_{j}\}_{j=1}^{\mathbf{N}}.$ In this section we elaborate on the method of Section 6 when $\mathbf{N}$ takes the values of $1,$ $2,$ and $3.$ We remark that in this section we no longer suppress the $t$-dependence in the relevant quantities.
When $\mathbf{N}=1,$ we use the input data set $\{\eta_{1},r_{1}\}$ to have the $1$-soliton solution $Q(x,t)$ via (6.17) by constructing the key quantity $\Delta(x,t)$ appearing in (6.22). In this case, (6.22) reduces to
$\Delta(x,t)=1+y_{1},$ (7.1)
where $y_{1}$ is obtained from (6.20) with $\chi_{1}$ being the the quantity we get from (6.21) when $j=1.$ The coefficient $\alpha_{11}/(2\alpha_{0})$ in (6.20) is evaluated with the help of (5.54), (6.14), and (6.18). By using (5.54) in (6.14), we obtain
$\frac{s_{1}}{r_{1}}=\begin{cases}\sqrt{3},\qquad P\equiv 0,\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr 0,\qquad P=Q_{x},\end{cases}$ (7.2)
where the first line refers to the case $P\equiv 0$ and the second line refers to the case $P=Q_{x}.$ Next, using (5.54) and (7.2) in (6.20), we obtain the quantity $y_{1}$ as
$y_{1}=\begin{cases}-2r_{1}\chi_{1},\qquad P\equiv 0,\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr r_{1}\chi_{1},\qquad P=Q_{x},\end{cases}$ (7.3)
where we recall that $\chi_{1}$ is the exponential function in (6.21) when $j=1.$ We remark that either of the two lines in (7.3) yields two equivalent values for $y_{1}$ by restricting the parameter $r_{1}$ in the first line to negative values and by restricting the parameter $y_{1}$ in the second line to positive values. Using (7.1) and (7.3) in (6.17), we obtain the $1$-soliton solution to (2.1) as
$Q(x,t)=\begin{cases}-\displaystyle\frac{36r_{1}\eta_{1}^{2}\chi_{1}}{\left(1-2r_{1}\chi_{1}\right)^{2}},\qquad P\equiv 0,\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr\displaystyle\frac{18r_{1}\eta_{1}^{2}\chi_{1}}{\left(1+r_{1}\chi_{1}\right)^{2}},\qquad P=Q_{x},\end{cases}$ (7.4)
where it is understood that $r_{1}<0$ in the first line and $r_{1}>0$ in the second line. We remark that the substitution $r_{1}\mapsto-r_{1}/2$ in the first line yields the second line in (7.4).
In Figure 7.1 we show the behavior of the $1$-soliton solution corresponding to the input data set with the parameter values given by
$(\eta_{1},r_{1})=\left(1,1\right),$ (7.5)
where we use the second line of (7.4) corresponding to $P=Q_{x}.$ By using the method of Section 6, we explicitly construct the $1$-soliton solution with the input data set in (7.5). The $1$-soliton solution $Q(x,t)$ in (7.4) is a real-valued solitary wave moving from the left to the right with the constant speed $9\eta_{1}^{4},$ which is equal to $9$ in this particular case. We remark that the $1$-soliton behavior for the Sawada–Kotera equation (2.1) illustrated in Figure 7.1 is similar to the behavior [1, 2, 11, 12, 23] of the $1$-soliton solution to the KdV equation in the sense that the soliton moves from the left to the right without changing its shape.
Figure 7.1: The snapshots for the $1$-soliton solution $Q(x,t)$ to (2.1) at t=0.5,t=-0.5, $t=0,$ and $t=0.5,$ respectively.
When $\mathbf{N}=2,$ we use the input data set $\{\eta_{1},\eta_{2},r_{1},r_{2}\}$ to construct the $2$-soliton solution $Q(x,t)$ to (2.1) via (6.17) by forming the quantity $\Delta(x,t)$ appearing in (6.22). In this case, (6.22) yields
$\Delta(x,t)=1+y_{1}+y_{2}+A_{12}\,y_{1}\,y_{2},$ (7.6)
where $A_{12}$ is obtained from (6.24) and the quantities $y_{1}$ and $y_{2}$ are constructed with the help of (6.15) and (6.18)–(6.21). Using (5.54) in (6.14) and exploiting the symmetrical appearance of $k_{1}$ and $k_{2}$ in (5.2), we obtain
$\frac{s_{1}}{r_{1}}=\begin{cases}\displaystyle\frac{\sqrt{3}\,\eta_{2}^{2}}{2\eta_{1}^{2}+\eta_{2}^{2}},\qquad P\equiv 0,\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr-\displaystyle\frac{\sqrt{3}\,\eta_{1}^{2}}{\eta_{1}^{2}+2\eta_{2}^{2}},\qquad P=Q_{x},\end{cases}\quad\displaystyle\frac{s_{2}}{r_{2}}=\begin{cases}\displaystyle\frac{\sqrt{3}\,\eta_{1}^{2}}{\eta_{1}^{2}+2\eta_{2}^{2}},\qquad P\equiv 0,\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr-\displaystyle\frac{\sqrt{3}\,\eta_{2}^{2}}{2\eta_{1}^{2}+\eta_{2}^{2}},\qquad P=Q_{x},\end{cases}$ (7.7)
where the first line refers to the case $P\equiv 0$ and the second line refers to the case $P=Q_{x}.$ Next, using (5.54) and (7.7) in (6.18), we obtain the quantity $y_{1}$ appearing in (6.20) and (7.6) as
$y_{1}=\begin{cases}\displaystyle\frac{2r_{1}\left(\eta_{1}^{3}+2\eta_{1}^{2}\eta_{2}+2\eta_{1}\eta_{2}^{2}+\eta_{2}^{3}\right)\chi_{1}}{\left(\eta_{1}-\eta_{2}\right)\left(2\eta_{1}^{2}+\eta_{2}^{2}\right)},\qquad P\equiv 0,\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr-\displaystyle\frac{2r_{1}\left(\eta_{1}^{3}+2\eta_{1}^{2}\eta_{2}+2\eta_{1}\eta_{2}^{2}+\eta_{2}^{3}\right)\chi_{1}}{\left(\eta_{1}-\eta_{2}\right)\left(\eta_{1}^{2}+2\eta_{2}^{2}\right)},\qquad P=Q_{x},\end{cases}$ (7.8)
with $\chi_{1}$ being the exponential function obtained from (6.21) when $j=1.$ As a result of the ordering $0<\eta_{1}<\eta_{2},$ in the first line of (7.8) we use $r_{1}<0$ and in the second line we use $r_{1}>0$ so that the quantity $y_{1}$ remains positive for all real values of $x$ and $t.$ The value of $y_{2}$ is obtained with the help of (7.8) by using the replacement $(\eta_{1},\eta_{2},r_{1})\mapsto(\eta_{2},\eta_{1},r_{2}).$ Thus, we get
$y_{2}=\begin{cases}-\displaystyle\frac{2r_{2}\left(\eta_{1}^{3}+2\eta_{1}^{2}\eta_{2}+2\eta_{1}\eta_{2}^{2}+\eta_{2}^{3}\right)\chi_{2}}{\left(\eta_{1}-\eta_{2}\right)\left(\eta_{1}^{2}+2\eta_{2}^{2}\right)},\qquad P\equiv 0,\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr\displaystyle\frac{2r_{2}\left(\eta_{1}^{3}+2\eta_{1}^{2}\eta_{2}+2\eta_{1}\eta_{2}^{2}+\eta_{2}^{3}\right)\chi_{2}}{\left(\eta_{1}-\eta_{2}\right)\left(2\eta_{1}^{2}+\eta_{2}^{2}\right)},\qquad P=Q_{x},\end{cases}$ (7.9)
with $\chi_{2}$ being the exponential function obtained from (6.21) when $j=2.$ In order to have $y_{2}$ to remain positive for all real values of $x$ and $t,$ as a result of the ordering $0<\eta_{1}<\eta_{2},$ we see that we must use $r_{2}>0$ in the first line of (7.9) and use $r_{2}<0$ in the second line.
In Figure 7.2 we show the behavior of the $2$-soliton solution corresponding to the input data set $\{\eta_{1},\eta_{2},r_{1},r_{2}\}$ with the specific values of the parameters given by
$(\eta_{1},\eta_{2},r_{1},r_{2})=\left(1,\displaystyle\frac{11}{10},1,-1\right),$ (7.10)
where we use the second lines in (7.7)–(7.9) corresponding to the case $P=Q_{x}.$ Using the method of Section 6, we explicitly construct the $2$-soliton solution with the input data set related to (7.10). The left plot in Figure 7.2 shows a snapshot when the two solitons are apart from each other and not yet interacting with each other. The next plot shows a snapshot at the beginning of the nonlinear interactions between the two solitons. The third plot shows a snapshot at the end of the nonlinear interactions. Finally, the right plot shows a snapshot when the two solitons are apart from each other and there are no longer any interactions. Both solitons travel from the left to the right. Initially, before the nonlinear interactions start, the taller soliton behind has a higher speed than the shorter soliton in the front. In this case, the initial speeds are $13.18$ and $9,$ respectively. We remark that the speed value $13.18$ corresponds to $9\eta_{2}^{4}$ and the speed value $9$ corresponds to $9\eta_{1}^{4}$ in this case. During the nonlinear interactions, the energy from the soliton behind is transmitted to the soliton in the front. Once the nonlinear interactions are over, the speed of the taller soliton behind is reduced from $13.18$ to $9$ and its height is reduced from $5.4$ to $4.5.$ On the other hand, the speed of the shorter soliton in the front has increased from $9$ to $13.18,$ and its height has increased from $2.9$ to $3.7.$ The soliton behind is still taller than the soliton in the front, but it is less taller after the nonlinear interactions. The $2$-soliton behavior for the Sawada–Kotera equation observed in this case is different from the $2$-soliton behavior [1, 2, 11, 12, 23] for the KdV equation. For the KdV equation, although both solitons move from the left to the right, each soliton retains its speed and its shape after the nonlinear interactions. For the KdV equation, the taller soliton always has a higher speed than the shorter soliton.
Figure 7.2: The snapshots for the $2$-soliton solution $Q(x,t)$ to (2.1) with the parameter values in (7.10) at t=2,t=-2, t=0.5,t=-0.5, $t=0,$ and $t=2,$ respectively.
In Figure 7.3 we illustrate the behavior of the $2$-soliton solution corresponding to the input data set $\{\eta_{1},\eta_{2},r_{1},r_{2}\}$ with the parameter values given by
$(\eta_{1},\eta_{2},r_{1},r_{2})=\left(\displaystyle\frac{17}{10},\displaystyle\frac{18}{10},1,-1\right),$ (7.11)
where we again use $P=Q_{x}.$ We explicitly construct the $2$-soliton solution with the input data set corresponding to (7.11). As in Figure 7.2, the left plot in Figure 7.3 shows a snapshot when the two solitons are apart from each other and not yet interacting with each other, the next plot shows a snapshot at the beginning of the nonlinear interactions, the third plot shows a snapshot at the end of the nonlinear interactions, and the right plot shows a snapshot when the two solitons are apart from each other and there are no longer any interactions. Again, both solitons travel from the left to the right. Initially, before the nonlinear interactions start, the shorter soliton behind has a higher speed than the taller soliton in the front, with the speeds being $94.48$ and $75.17,$ respectively. We remark that the speed value $94.48$ corresponds to $9\eta_{2}^{4}$ and the speed value $75.17$ corresponds to $9\eta_{1}^{4}.$ During the nonlinear interactions, the energy from the soliton behind is transmitted to the soliton in the front. Once the nonlinear interactions are over, the speed of the shorter soliton behind is reduced from $94.48$ to $75.17$ and its height is reduced from $14.6$ to $13.$ On the other hand, the speed of the taller soliton in the front has increased from $75.17$ to $94.48,$ and its height has increased from $55.1$ to $58.3.$ The soliton behind is still shorter than the soliton in the front. The $2$-soliton behavior for the Sawada–Kotera equation observed in this case is again different from the $2$-soliton behavior for the KdV equation, with the exception of the fact that for both equations all the solitons move from the left to the right.
Figure 7.3: The snapshots for the $2$-soliton solution $Q(x,t)$ to (2.1) with the parameter values in (7.11) at t=0.2,t=-0.2, t=0.08,t=-0.08, $t=0,$ and $t=0.2,$ respectively.
When $\mathbf{N}=3,$ by using the input data set $\{\eta_{1},\eta_{2},\eta_{3},r_{1},r_{2},r_{3}\},$ we construct the $3$-soliton solution to (2.1). As explained in Section 6, we first construct the quantity $\Delta(x,t)$ given in (6.22) and then obtain the $3$-soliton solution $Q(x,t)$ via (6.17). When $\mathbf{N}=3,$ from (6.22) we get
$\Delta(x,t)=1+y_{1}+y_{2}+y_{3}+A_{12}\,y_{1}\,y_{2}+A_{13}\,y_{1}\,y_{3}+A_{23}\,y_{2}\,y_{3}+A_{12}\,A_{13}\,A_{23}\,y_{1}\,y_{2}\,y_{3},$ (7.12)
where the coefficients $A_{12},$ $A_{13},$ $A_{23},$ are obtained in terms of $\eta_{1},$ $\eta_{2},$ $\eta_{3}$ by using (6.24). We have
$A_{12}=\displaystyle\frac{(\eta_{1}-\eta_{2})^{3}(\eta_{1}^{3}+\eta_{2}^{3})}{(\eta_{1}+\eta_{2})^{3}(\eta_{1}^{3}-\eta_{2}^{3})},\quad A_{13}=\displaystyle\frac{(\eta_{1}-\eta_{3})^{3}(\eta_{1}^{3}+\eta_{3}^{3})}{(\eta_{1}+\eta_{3})^{3}(\eta_{1}^{3}-\eta_{3}^{3})},\quad A_{23}=\displaystyle\frac{(\eta_{2}-\eta_{3})^{3}(\eta_{2}^{3}+\eta_{3}^{3})}{(\eta_{2}+\eta_{3})^{3}(\eta_{2}^{3}-\eta_{3}^{3})}.$
We remark that the values $A_{13}$ and $A_{23}$ can readily be obtained from the value $A_{12}$ by exploiting the symmetrical appearance of $k_{1},$ $k_{2},$ and $k_{3}$ in (5.2). The quantities $y_{1},$ $y_{2},$ $y_{3}$ are as in (6.20) with $\chi_{1},$ $\chi_{2},$ $\chi_{3}$ being the exponential functions appearing in (6.21). Thus, we have
$y_{1}=\displaystyle\frac{\alpha_{11}}{2\alpha_{0}}\,\chi_{1},\quad y_{2}=\displaystyle\frac{\alpha_{22}}{2\alpha_{0}}\,\chi_{2},\quad y_{3}=\displaystyle\frac{\alpha_{33}}{2\alpha_{0}}\,\chi_{3},$
where the three coefficients $\alpha_{jj}/(2\alpha_{0})$ for $j=1,2,3$ are obtained with the help of (6.19). We then construct the $3$-soliton solution $Q(x,t)$ by using (7.12) in (6.17).
In Figure 7.4 we illustrate the behavior of the $3$-soliton solution corresponding to the input data set $\{\eta_{1},\eta_{2},\eta_{3},r_{1},r_{2},r_{3}\}$ with the parameter values given by
$(\eta_{1},\eta_{2},\eta_{3},r_{1},r_{2},r_{3})=\left(\displaystyle\frac{32}{100},\displaystyle\frac{36}{100},\displaystyle\frac{4}{10},8,-8,8\right),$ (7.13)
where we use the case $P=Q_{x}.$ Via the method of Section 6, we explicitly construct the $3$-soliton solution with the input data set corresponding to (7.13). The left plot in Figure 7.4 shows a snapshot when the three solitons are apart from each other and not yet interacted with each other. The next plot shows a snapshot at the beginning of the nonlinear interactions among the three solitons. The third plot shows a snapshot at the end of the nonlinear interactions. Finally, the right plot shows a snapshot when the three solitons are apart from each other and they no longer interact. All three solitons travel from the left to the right. Initially, before the nonlinear interactions start, the tallest soliton behind has a higher speed than the two shorter solitons in the front. In this case, the initial speeds are $0.23,$ $0.151,$ and $0.094,$ respectively, for the tallest soliton behind, the middle-height soliton in the middle, and the shortest soliton in the front. We remark that the speed value $0.23$ corresponds to $9\eta_{3}^{4},$ the speed value $0.151$ corresponds to $9\eta_{2}^{4},$ and the speed value $0.094$ corresponds to $9\eta_{1}^{4}.$ During the nonlinear interactions, the energy from each soliton behind is transmitted to the next soliton in the front. Once the nonlinear interactions are over, the three solitons are again aligned so that the tallest soliton is still behind, the middle-height soliton is still in the middle, and the shortest soliton is still in the front. However, the speed of the tallest soliton has decreased from $0.23$ to $0.094,$ the speed of the middle-height soliton remains unchanged at the speed value $0.151,$ and the speed of the shortest soliton has increased from $0.094$ to $0.23.$ On the other hand, the height of the tallest soliton behind has decreased from $0.72$ to $0.46,$ the height of the middle-height soliton in the middle has decreased from $0.1$ to $0.065,$ and the height of the shortest soliton in the front has decreased from $0.055$ to $0.017.$ The $3$-soliton behavior for the Sawada–Kotera equation observed in this case is different from the $3$-soliton behavior for the KdV equation. For the KdV equation, although all three solitons move from the left to the right, each soliton retains its speed and its shape after the nonlinear interactions, and a taller soliton always has a higher speed than a shorter soliton.
Figure 7.4: The snapshots for the $3$-soliton solution $Q(x,t)$ to (2.1) with the parameter values in (7.13) at t=300,t=-300, t=100,t=-100, $t=0,$ and $t=300,$ respectively.
In Figure 7.5 we illustrate the behavior of the $3$-soliton solution corresponding to the input data set $\{\eta_{1},\eta_{2},\eta_{3},r_{1},r_{2},r_{3}\}$ with the parameter values given by
$(\eta_{1},\eta_{2},\eta_{3},r_{1},r_{2},r_{3})=\left(1,\displaystyle\frac{11}{10},\displaystyle\frac{12}{10},\displaystyle\frac{1}{10},-\displaystyle\frac{1}{20},1\right),$ (7.14)
where we again use the case $P=Q_{x}.$ We explicitly construct the $3$-soliton solution with the input data set corresponding to (7.14). The left plot in Figure 7.5 shows a snapshot when the three solitons are apart from each other and not yet interacting with each other. The next plot shows a snapshot at the beginning of the nonlinear interactions. The third plot shows a snapshot at the end of the nonlinear interactions. Finally, the right plot shows a snapshot when the three solitons are apart from each other and they no longer interact. All three solitons travel from the left to the right. Initially, before the nonlinear interactions start, the tallest soliton in the front has a lower speed than the middle-height soliton behind and the shortest soliton in the middle. In this case, the initial speeds are $18.66,$ $13.18,$ and $9,$ respectively, for the middle-height soliton behind, the shortest soliton in the middle, and the tallest soliton in the front. During the nonlinear interactions, the energy from each soliton behind is transmitted to the soliton in the front. Once the nonlinear interactions are over, the three solitons are again aligned so that the tallest soliton is still in the front, the shortest soliton is still in the middle, and the middle-height soliton is still behind. However, the speed of the tallest soliton in the front has increased from $9$ to $18.66,$ the speed of the shortest soliton in the middle has remained unchanged at the value $13.18,$ and the speed of the middle-height soliton behind has decreased from $18.66$ to $9.$ On the other hand, the height of the tallest soliton in the front has increased from $12.6$ to $19,$ the height of the shortest soliton in the middle has decreased from $5.3$ to $3.7,$ and the height of the middle-height soliton behind has decreased from $6.5$ to $4.5.$ The behavior of the $3$-soliton solution to the Sawada–Kotera equation observed in this case is different from the $3$-soliton behavior for the KdV equation. For the KdV equation, although all three solitons move from the left to the right, each soliton retains its speed and its shape after the nonlinear interactions, and a taller soliton always has a higher speed than a shorter soliton.
Figure 7.5: The snapshots for the $3$-soliton solution $Q(x,t)$ to (2.1) with the parameter values in (7.14) at t=2,t=-2, t=0.5,t=-0.5, $t=0,$ and $t=2,$ respectively.